### v1

Abstract
This paper studies the implementation of Bayes correlated equilibria in symmetric Bayesian nonatomic games, using direct information structures and obedient strategies. The main results demonstrate full implementation in a class of games with positive cost externalities. Specifically, if the game admits a strictly convex potential in every state, then for every Bayes correlated equilibrium outcome with finite support and rational action distributions, there exists a direct information structure that implements this outcome under all equilibria. When the potential is only weakly convex, we show that all equilibria implement the same expected social cost. Additionally, all Bayes correlated equilibria, including those with infinite support or irrational action distributions, are approximately implemented.
Keywords: Bayes correlated equilibrium, Bayesian Wardrop equilibrium, congestion games, information design, nonatomic games, potential games.
1 Introduction
In this paper, we study partial and full implementation through information design in symmetric cost games with incomplete information and a continuum of a priori uninformed players. A perfectly informed designer can choose how to correlate players’ information with the payoff-relevant state to influence their actions, for example, to minimize a cost function such as the social cost. In this context, the concept of Bayes correlated Wardrop equilibrium (BCWE) (the analog of a Bayes correlated equilibrium in games with a finite number of players) characterizes the set of state-dependent probability distributions over action distributions (hereafter called flows) that can be induced in an incentive-compatible way by some information structure. [^4] Koessler et al., (2024) show that the set of BCWE is the natural limit of the set of Bayes correlated equilibria (Bergemann and Morris,, 2016) in finite games with a large number of players. Furthermore, as shown in Koessler et al., (2024), regardless of the designer’s objective, a designer-optimal BCWE can be formulated as a solution to an optimization program with a bounded number of variables and constraints, determined by the number of actions and states.
We show that approaching information design through the lens of nonatomic games provides important insights, particularly for achieving full implementation through direct recommendation systems. There is full implementation when all equilibria implement the same outcome for the chosen information structure. This property is crucial because it ensures the robustness of the information design problem to adversarial equilibrium selection. When only some equilibria induce the desired outcome, we talk about partial implementation.
Our basic game is a nonatomic game with symmetric and incomplete information, thereby extending the class of population games studied by Sandholm, (2001, 2010) to the Bayesian setting. [^5] Sandholm, (2001, 2010) allow for multiple populations that differ in their preferences and sets of feasible actions. For notational simplicity, we assume in this paper that the basic game involves only one population, so players are ex-ante symmetric. We expect our results to hold more generally for multiple populations. We extend this basic game to asymmetric information by considering a family of simple information structures (as defined in Wu et al.,, 2021): The set of players is partitioned into a finite number of populations; within any given population, all players receive the same signal. Such an information structure is a direct information structure if the set of possible signals is identical to the set of possible actions. A Bayesian Wardrop equilibrium (BWE) (the analog of Bayes-Nash equilibrium for games with finite number of players) of the game extended with an information structure is said to be obedient if it is such that all players in all populations follow their recommended action.
We first show that every BCWE with finite support and rational flows can be partially implemented as an obedient BWE with an appropriate direct information structure (Theorem 4.1). Hence, a version of the revelation principle (Myerson,, 1982) applies to symmetric nonatomic Bayesian games. Next, we turn to full implementation, which ensures that every BWE outcome of the extended game leads to the desired BCWE. We show that if the basic game admits a strictly convex potential in each state, then the BWE outcome is unique for every information structure (Theorem 5.3). Consequently, information design becomes extremely powerful in nonatomic games with a strictly convex potential because every BCWE with finite support and rational flows can be fully implemented—there exists a direct information structure such that the BCWE is implemented, regardless of the BWE selection.
If in each state the basic game admits a potential that is only weakly convex, full implementation of a given outcome is usually not feasible via information design. Yet, we show that if the potential is weakly convex, then for every BCWE with finite support and rational flows, there exists a direct information structure such that the expected social cost is the same as the BCWE expected social cost regardless of the BWE (Theorem 5.7). In particular, the expected cost of the social cost-minimizing BCWE can be fully implemented.
[…]
Our full implementation results apply, in particular, to congestion games with increasing cost functions. A congestion game has a strictly convex potential if the set of actions is equal to the set of resources and the cost of using a resource is strictly increasing in the number of users. When actions are general feasible subsets of resources and costs are added across resources, congestion games admit a weakly convex potential, provided that the cost of using each resource is weakly increasing in the number of users.
1.1 Related Literature
The notion of Wardrop equilibrium originated from the work of Wardrop, (1952) on traffic flow theory, where he formulated principles describing user-optimal and system-optimal routing in transportation networks. Beckmann et al., (1956) proved that a Wardrop equilibrium is the solution of a convex optimization problem. This idea was framed in the language of potential games (see, among others, Sandholm,, 2001, Roughgarden,, 2007, Roughgarden and Tardos,, 2007), using a generalization of the concept of potential game from the case of a finite number of players (Rosenthal,, 1973, Monderer and Shapley,, 1996) to the nonatomic case. A general framework for nonatomic potential games, extending beyond routing games, was studied by Sandholm, (2001), Hofbauer and Sandholm, (2009), Sandholm, (2010). Koessler et al., (2024) extend this framework to Bayesian games with nonatomic players by studying the notions of correlated and Bayes correlated Wardrop equilibrium. They showed that Bayes correlated Wardrop equilibria emerge as limits of action flows induced by Bayes correlated equilibria (Bergemann and Morris,, 2016) in finite-player games. Related nonatomic versions of correlated and Bayes correlated equilibria were studied by, e.g., Díaz et al., (2009), Tavafoghi and Teneketzis, (2018), and Zhu and Savla, (2022). Acemoglu et al., (2018) showed the relevance of information design in routing games, and proved that some classical paradoxes in routing games can be expressed in terms of negative value of information. Among the recent contributions on the topic, the reader is referred, for instance, to Massicot and Langbort, (2022), Cianfanelli et al., (2023), Ambrogio et al., (2024), Griesbach et al., (2024).
To the best of our knowledge, our full implementation results with direct information structures have no analog in the literature on information design, where even for binary action games, full implementation typically relies on complex indirect information structures, often with infinite type spaces. Existing tools and recent results in this area include, for example, Mathevet et al., (2020), Dworczak and Pavan, (2022), Hoshino, (2022), Inostroza and Pavan, (2023), Li et al., (2023), and Morris et al., (2024). The literature on information design is surveyed in Bergemann and Morris, (2019), Kamenica, (2019) and Forges, (2020). Dughmi, (2017) surveys the algorithmic literature on information design.
1.2 Organization of the Paper
The next section presents the model, including the basic game, its Bayes correlated Wardrop equilibria, information structures with finite populations and public signals within populations, and Bayesian Wardrop ($\varepsilon$-)equilibria of the basic game extended with an information structure. Section 3 provides two simple examples and illustrates how to implement some BCWE. Our partial and full implementation results are presented in Sections 4 and 5, respectively.
1.3 Notation
Given any finite set $A$, we define
$\Delta_{\gamma}(A)\coloneqq\left\{\boldsymbol{y}=\left(y_{a}\right)_{a}\colon y_{a}\geq 0,\sum_{a\in A}y_{a}=\gamma\right\}.$ (1.1)
For the sake of simplicity, we use the shorthand $\Delta(A)$ to indicate $\Delta_{1}(A)$, which is just the simplex of probability measures on $A$. For a compact set $\mathcal{Y}$, $\Delta{(\mathcal{Y})}$ denotes the set of Borel probability measures on $\mathcal{Y}$.
2 The Model
We consider a basic game $\Gamma$, which is a symmetric Bayesian game with nonatomic players, a finite set of actions, and a finite set of states. Following Koessler et al., (2024) and most of the literature on nonatomic games, we adopt the convention that players minimize cost functions, which are simply the negatives of utility functions. Formally:
Definition 2.1.
A basic game $\Gamma=\left(\mathcal{A},\Theta,p,\boldsymbol{c}\right)$ is given by a finite set of actions $\mathcal{A}$, a finite set of states $\Theta$, a full-support probability distribution over states $p\in\Delta(\Theta)$, and a profile of cost functions $\boldsymbol{c}=(c_{a})_{a\in\mathcal{A}}$, where for each $a\in\mathcal{A}$, $c_{a}\colon\mathcal{Y}\times\Theta\to\mathbb{R}$ is a continuous cost function and $\mathcal{Y}=\Delta(\mathcal{A})$ is the set of flow profiles.
A flow profile (or action distribution) is denoted by $\boldsymbol{y}=(y_{a})_{a\in\mathcal{A}}\in\mathcal{Y}$, where $y_{a}$ is the flow on action $a$. For every $a\in\mathcal{A}$, $\theta\in\Theta$ and $\boldsymbol{y}\in\mathcal{Y}$, $c_{a}(\boldsymbol{y},\theta)$ is the cost of action $a$ when the state is $\theta$ and the flow profile in that state is $\boldsymbol{y}$. A transition probability $\mu\colon\Theta\to\Delta(\mathcal{Y})$ that associates a distribution over flows to each state is called an outcome of the game.
The notion of Bayes correlated Wardrop equilibrium in the next definition is the analog of the notion of Bayes correlated equilibrium (Bergemann and Morris,, 2016) in the context of nonatomic games (see Koessler et al.,, 2024).
Definition 2.2.
A Bayes correlated Wardrop equilibrium (BCWE) of $\Gamma$ is an outcome $\mu:\Theta\to\Delta(\mathcal{Y})$ such that
$\sum_{\theta\in\Theta}p(\theta)\int y_{a}c_{a}(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)\leq\sum_{\theta\in\Theta}p(\theta)\int y_{a}c_{b}(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta),\quad\forall\,a,b\in\mathcal{A}.$ (2.1)
In words, for each state $\theta\in\Theta$, a flow profile $\boldsymbol{y}$ is drawn at random according to the distribution $\mu(\,\cdot\mid\theta)$. The corresponding outcome is a BCWE if, for every action $a$ that has strictly positive flow in at least one state, conditional on being assigned $a$, the expected cost of playing $a$ is not strictly higher than the expected cost of playing any other action $b$. For example, suppose that for every $\theta$, $\boldsymbol{y}(\theta)$ is a Wardrop equilibrium of the complete information game $(\mathcal{A},(c_{a}(\,\cdot\,,\theta))_{a\in\mathcal{A}})$ i.e.,
$\forall a,b\in\mathcal{A},\ y_{a}(\theta)>0\implies c_{a}(\boldsymbol{y}(\theta),\theta)\leq c_{b}(\boldsymbol{y}(\theta),\theta).$ (2.2)
Then $\mu(\theta)=\delta_{\boldsymbol{y}(\theta)}$ is a “fully revealing” BCWE.
As another example, if $\bar{\boldsymbol{y}}$ is a Wardrop equilibrium of the average game with complete information $(\mathcal{A},(\bar{c}_{a})_{a\in\mathcal{A}})$, with $\bar{c}_{a}(\,\cdot\,)=\sum_{\theta}p(\theta)c_{a}(\,\cdot\,,\theta)$, then $\mu(\theta)=\delta_{\bar{\boldsymbol{y}}}$ is a “non-revealing” BCWE. This corresponds to an equilibrium outcome of the basic game with no information.
We now introduce finite information structures and Bayesian Wardrop equilibria as defined in Wu et al., (2021).
Definition 2.3.
An information structure $\mathcal{I}=\left(\boldsymbol{\gamma},\mathcal{T},\pi\right)$ for $\Gamma$ is given by:
  • a vector $\boldsymbol{\gamma}=(\gamma^{k})_{k\in\mathcal{K}}\in\Delta(\mathcal{K})$, where $\mathcal{K}$ is a finite set of populations and, for every $k\in\mathcal{K}$, $\gamma^{k}>0$ is the size of population $k$;
  • for each population $k$, a finite set of types $\mathcal{T}^{k}$;
  • a mapping $\pi\colon\Theta\to\Delta(\mathcal{T})$, where $\mathcal{T}=\times_{k\in\mathcal{K}}\mathcal{T}^{k}$.
The game $\Gamma$ extended by information structure $\mathcal{I}$, denoted $(\Gamma,\mathcal{I})$, can be described as follows: for each state $\theta\in\Theta$, a type profile $\boldsymbol{\tau}\in\mathcal{T}$ is drawn with probability $\pi(\boldsymbol{\tau}\mid\theta)$ and, for every $k$, the entire population $k$ observes the type $\tau^{k}\in\mathcal{T}^{k}$. Then, actions are chosen.
For all $k\in\mathcal{K}$, $\tau^{k}\in\mathcal{T}^{k}$ and $a\in\mathcal{A}$, we let $y_{a}^{k}(\tau^{k})$ denote the flow on action $a$ of population $k$, when type $\tau^{k}$ is observed. The interim flow profile of population $k$ given type $\tau^{k}$ is $\boldsymbol{y}^{k}(\tau^{k})=\left(y_{a}^{k}(\tau^{k})\right)_{a\in\mathcal{A}}\in\Delta_{\gamma^{k}}(\mathcal{A})$. The list
$\widehat{\boldsymbol{y}}=\left(\boldsymbol{y}^{k}(\tau^{k})\right)_{k,\tau^{k}}\in\bigtimes_{k\in\mathcal{K}}\left[\bigtimes_{\tau^{k}\in\mathcal{T}^{k}}\Delta_{\gamma^{k}}(\mathcal{A})\right],$ (2.3)
of interim flows profiles of all populations is called interim flow profile. For each profile of types $\boldsymbol{\tau}=(\tau^{k})_{k\in\mathcal{K}}$, we let $y_{a}(\boldsymbol{\tau})=\sum_{k\in\mathcal{K}}y_{a}^{k}(\tau^{k})$ be the interim total flow on action $a$ and $\boldsymbol{y}(\boldsymbol{\tau})=\left(y_{a}(\boldsymbol{\tau})\right)_{a\in\mathcal{A}}\in\Delta(\mathcal{A})$ be the interim total flow profile given $\boldsymbol{\tau}$. We let $\operatorname{\mathsf{P}}(\tau^{k})=\sum_{\theta\in\Theta}\sum_{\boldsymbol{\tau}^{-k}\in\mathcal{T}^{-k}}p(\theta)\pi(\tau^{k},\boldsymbol{\tau}^{-k}\mid\theta)$ denote the total probability of type $\tau^{k}$ induced by $p$ and $\pi$, where $\mathcal{T}^{-k}\coloneqq\times_{j\neq k}\mathcal{T}^{j}$. Under the interim flow profile $\widehat{\boldsymbol{y}}$, the cost of population $k$, given type $\tau^{k}$ and action $a$ such that $y_{a}^{k}(\tau^{k})>0$, is
$\sum_{\theta\in\Theta}\sum_{\boldsymbol{\tau}^{-k}\in\mathcal{T}^{-k}}\frac{p(\theta)\pi(\tau^{k},\boldsymbol{\tau}^{-k}\mid\theta)}{\operatorname{\mathsf{P}}(\tau^{k})}c_{a}(\boldsymbol{y}(\boldsymbol{\tau}),\theta)\eqqcolon c_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}}).$ (2.4)
The list $(c_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}}))_{k,\tau^{k},a}$ is called an interim cost profile.
The following definition provides an analog of the usual Bayesian Nash ($\varepsilon$-)equilibrium for nonatomic games.
Definition 2.4.
[…]
$\sum_{\theta\in\Theta}\sum_{\boldsymbol{\tau}^{-k}\in\mathcal{T}^{-k}}\frac{p(\theta)\pi(\tau^{k},\boldsymbol{\tau}^{-k}\mid\theta)}{\operatorname{\mathsf{P}}(\tau^{k})}c_{a}(\boldsymbol{y}(\boldsymbol{\tau}),\theta)\leq\sum_{\theta\in\Theta}\sum_{\boldsymbol{\tau}^{-k}\in\mathcal{T}^{-k}}\frac{p(\theta)\pi(\tau^{k},\boldsymbol{\tau}^{-k}\mid\theta)}{\operatorname{\mathsf{P}}(\tau^{k})}c_{b}(\boldsymbol{y}(\boldsymbol{\tau}),\theta)+\varepsilon.$ (2.5)
That is, for each population $k$ that observes type $\tau^{k}$, a positive mass chooses action $a$ only if, up to $\varepsilon$, it has the smallest conditional expected cost. Notice that this is equivalent to the following inequality for all $k\in\mathcal{K}$, all $\tau^{k}\in\mathcal{T}^{k}$ and all $a,b\in\mathcal{A}$ such that $y_{a}^{k}(\tau^{k})>0$:
$\sum_{\theta\in\Theta}\sum_{\boldsymbol{\tau}^{-k}\in\mathcal{T}^{-k}}p(\theta)\pi(\tau^{k},\boldsymbol{\tau}^{-k}\mid\theta)c_{a}(\boldsymbol{y}(\boldsymbol{\tau}),\theta)\leq\sum_{\theta\in\Theta}\sum_{\boldsymbol{\tau}^{-k}\in\mathcal{T}^{-k}}p(\theta)\pi(\tau^{k},\boldsymbol{\tau}^{-k}\mid\theta)c_{b}(\boldsymbol{y}(\boldsymbol{\tau}),\theta)+\varepsilon\operatorname{\mathsf{P}}(\tau^{k}).$ (2.6)
A Bayesian Wardrop equilibrium (BWE) is a Bayesian Wardrop $\varepsilon$-equilibrium with $\varepsilon=0$.
An outcome $\mu\colon\Theta\to\Delta(\mathcal{Y})$ is a Bayesian Wardrop $\varepsilon$-equilibrium outcome of $(\Gamma,\mathcal{I})$ if there exists a Bayesian Wardrop $\varepsilon$-equilibrium $\left(\boldsymbol{y}^{k}(\tau^{k})\right)_{k,\tau^{k}}$ such that for every $\theta\in\Theta$ and for every $\boldsymbol{z}\in\mathcal{Y}$,
$\mu(\boldsymbol{z}\mid\theta)=\sum_{\boldsymbol{\tau}\colon\boldsymbol{y}(\boldsymbol{\tau})=\boldsymbol{z}}\pi(\boldsymbol{\tau}\mid\theta).$ (2.7)
We let $\operatorname{\mathsf{BWE_{\varepsilon}}}(\mathcal{I})$ be the set of Bayesian Wardrop $\varepsilon$-equilibrium outcomes of $(\Gamma,\mathcal{I})$.
An information structure is direct if $\mathcal{T}^{k}=\mathcal{A}$ for all $k\in\mathcal{K}$. The associated obedient interim flow profile is the one obtained when population $k$ observes $\tau^{k}=a$, interprets it as an action recommendation, and obeys it by choosing $a$. For every direct information structure $\mathcal{I}$, we let $\breve{\boldsymbol{y}}_{\mathcal{I}}$ denote the obedient interim flow profile, defined as $\breve{\boldsymbol{y}}_{\mathcal{I},a}^{k}(\tau^{k})\coloneqq\gamma^{k}\cdot\operatorname{\mathds{1}}\left\{\tau^{k}=a\right\}$, for all $k\in\mathcal{K},\tau^{k}\in\mathcal{T}^{k},a\in\mathcal{A}$. The associated outcome is denoted by $\breve{\mu}_{\mathcal{I}}$.
3 Examples
The following examples illustrate the notion of BCWE in two basic games and show how to implement the social cost-minimizing BCWE as an obedient BWE of the basic game extended with a direct information structure. They also show that restricting attention to public information or deterministic flows is not without loss of generality.
Example 1.
[…]
$c_{a}(\boldsymbol{y},\theta)=\operatorname{\mathds{1}}\{\theta=\bar{\theta}\},\quad c_{b}(\boldsymbol{y},\theta)=2y_{b}+1/3.$ (3.1)
Assume that the designer aims at minimizing the social cost, defined as
$\operatorname{\mathsf{SC}}(\boldsymbol{y},\theta)=y_{a}c_{a}(\boldsymbol{y},\theta)+y_{b}c_{b}(\boldsymbol{y},\theta).$
[…]
This outcome can be implemented as a BWE with the following information structure: there are six populations of size $\gamma^{k}=1/6$ for $k=1,\ldots,6$, with
[…]
and
$\pi(b,a,a,a,a,a\mid\bar{\theta})=\pi(a,b,a,a,a,a\mid\bar{\theta})=\pi(a,a,b,a,a,a\mid\bar{\theta})=$
$\pi(a,a,a,b,a,a\mid\bar{\theta})=\pi(a,a,a,a,b,a\mid\bar{\theta})=\pi(a,a,a,a,a,b\mid\bar{\theta})=$ $1/6$
It is immediate to observe that the equilibrium conditions of the BWE are equivalent to those of the corresponding BCWE.
Example 2.
In the following basic game from Díaz et al., (2009) and Mitsche et al., (2013), there is a single state, and the action set is $\mathcal{A}=\left\{a,b\right\}$. The cost functions are:
$c_{a}(\boldsymbol{y})=1,\quad c_{b}(\boldsymbol{y})=\max\left\{2-4y_{b},4y_{b}-2\right\}.$ (3.3)
Action $b$ generates negative cost externalities when $y_{b}<1/2$ and positive cost externalities when $y_{b}>1/2$. This basic game without information admits three BWE outcomes: $(1,0)$, $(3/4,1/4)$, and $(1/4,3/4)$, all of which yield a social cost of $1$. Public information just induces a mixing over these outcomes. Now consider the flows $\boldsymbol{z}=(1,0)$ and $\boldsymbol{w}=(1/2,1/2)$. The outcome $\mu$, defined by
$\mu(\boldsymbol{z})=\frac{1}{3},\quad\mu(\boldsymbol{w})=\frac{2}{3},$ (3.4)
is a BCWE, since it satisfies the equilibrium constraints:
$\frac{2}{3}=\frac{1}{3}z_{a}c_{a}(\boldsymbol{z})+\frac{2}{3}w_{a}c_{a}(\boldsymbol{w})$ $\leq\frac{1}{3}z_{a}c_{b}(\boldsymbol{z})+\frac{2}{3}w_{a}c_{b}(\boldsymbol{w})=\frac{2}{3},$
$0=\frac{1}{3}z_{b}c_{b}(\boldsymbol{z})+\frac{2}{3}w_{b}c_{b}(\boldsymbol{w})$ $\leq\frac{1}{3}z_{b}c_{a}(\boldsymbol{z})+\frac{2}{3}w_{b}c_{a}(\boldsymbol{w})=\frac{1}{3}.$
The social cost of this BCWE is $2/3$, which is lower than the social cost with public information. The cost-minimizing BCWE is (partially) implemented as a BWE with the following direct information structure: there are two populations of size $\gamma^{1}=\gamma^{2}=1/2$, and
$\pi(a,a)=\pi(a,b)=\pi(b,a)=\frac{1}{3}.$
4 Partial Implementation
It is immediate to show that the set of all BWE for all possible (direct or indirect) information structures, as described in the previous section, is included in the set of BCWE. This simple observation explicitly illustrates how to construct some BCWE: choose an information structure $\left(\boldsymbol{\gamma},\mathcal{T},\pi\right)$ used to disclose information, and choose a BWE of the basic game extended with this information structure.
The next theorem establishes, at least approximately, the other direction of the revelation principle (Myerson,, 1982, Bergemann and Morris,, 2016). Specifically, it shows that if $\mu$ is a BCWE with finite support and rational flows, then there exists a direct information structure such that $\mu$ is implemented as an obedient BWE outcome. If the support of $\mu$ is not finite or if some flows in the support are not rational, then, for every $\varepsilon>0$, $\mu$ can be approximately implemented as a Bayesian Wardrop $\varepsilon$-equilibrium outcome.
Theorem 4.1.
Let $\mu$ be a BCWE and $\psi\colon\mathcal{Y}\times\Theta\to\mathbb{R}$ be a continuous function. For all $\varepsilon>0$, there exists a direct information structure $\mathcal{I}$ such that the obedient interim flow profile $\breve{\boldsymbol{y}}_{\mathcal{I}}$ is a Bayesian Wardrop $\varepsilon$-equilibrium of $(\Gamma,\mathcal{I})$ and,
$\left\lvert\int\psi(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)-\int\psi(\boldsymbol{y},\theta)\ \textup{d}\breve{\mu}_{\mathcal{I}}(\boldsymbol{y}\mid\theta)\right\rvert\leq\varepsilon,\quad\forall\theta\in\Theta.$ (4.1)
Further, if $\mu$ has finite support and if all flows in the support of $\mu$ are rational, then there exists an information structure $\mathcal{I}$ such that $\varepsilon=0$ and $\mu=\breve{\mu}_{\mathcal{I}}$.
In Theorem 4.1, $\psi$ represents any possible cost function for the designer, and (4.1) means that, for every state, the interim expected cost of the designer under outcome $\breve{\mu}_{\mathcal{I}}$ is $\varepsilon$-close to the interim expected cost under the BCWE $\mu$. The proof of the theorem also shows that, for every $L>0$, the same $\mathcal{I}$ can be chosen for all $L$-Lipschitz $\psi$.
Proof.
Koessler et al., (2024, Lemma 2) show that BCWE with finite support are dense within the set of Bayes correlated Wardrop equilibria. Therefore, it is enough to prove the result for $\mu$ with finite support. Fix $\varepsilon>0$ and a BCWE $\mu$ with finite support $\mathcal{Y}^{\ast}\subset\mathcal{Y}$. Let
$\varepsilon_{0}=\min\left\{y_{a}\text{ s.t. }\boldsymbol{y}\in\mathcal{Y}^{\ast},a\in\mathcal{A},y_{a}>0\right\}>0$ (4.2)
be the smallest positive flow in the finite set $\mathcal{Y}^{\ast}$.
We take an arbitrary $\eta>0$ and approximate flows in $\mathcal{Y}^{\ast}$ with vectors having rational components as follows. There exists $K\in\mathbb{N}_{+}$, such that for every $\boldsymbol{y}\in\mathcal{Y}^{\ast}$ and every $a\in\mathcal{A}$, there exists $N(y_{a})\in\mathbb{N}$, such that,
$\sum_{a\in\mathcal{A}}N(y_{a})=K\quad\text{and}\quad\forall a\in\mathcal{A},\left\lvert\frac{N(y_{a})}{K}-y_{a}\right\rvert\leq\eta.$ (4.3)
For each $\boldsymbol{y}\in\mathcal{Y}^{\ast}$, let $\tilde{\boldsymbol{y}}$ denote the flow profile $(\tilde{y}_{a})_{a\in\mathcal{A}}$ with $\tilde{y}_{a}=N(y_{a})/K$.
We construct a direct information structure $\mathcal{I}$ as follows. There are $|\mathcal{K}|=K$ populations and each population $k$ has the same size $\gamma^{k}=1/K$. For each population $k$, the set of types is $\mathcal{T}^{k}=\mathcal{A}$. Conditionally on state $\theta$, a flow profile $\boldsymbol{y}\in\mathcal{Y}^{\ast}$ is drawn with probability $\mu(\boldsymbol{y}\mid\theta)$, then action $a$ is recommended to a subset of $N(y_{a})$ populations chosen uniformly, so that, conditionally on $\boldsymbol{y}$ and $\theta$, the probability that population $k$ is recommended $a$ is $\tilde{y}_{a}$. We let $\breve{\mu}_{\mathcal{I}}$ be the induced obedient outcome which puts probability $\mu(\boldsymbol{y}\mid\theta)$ on $\tilde{\boldsymbol{y}}$ in state $\theta$.
For each population $k$, the conditional expected cost of playing $b$ when $a$ is recommended is
$\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}\frac{p(\theta)\mu(\boldsymbol{y}\mid\theta)\tilde{y}_{a}}{\operatorname{\mathsf{Q}}(a)}c_{b}(\tilde{\boldsymbol{y}},\theta),$ (4.4)
with
$\operatorname{\mathsf{Q}}(a)=\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}p(\theta)\mu(\boldsymbol{y}\mid\theta)\tilde{y}_{a}.$ (4.5)
For each pair of actions $(a,b)$, the mapping $\boldsymbol{y}\mapsto y_{a}c_{b}(\boldsymbol{y},\theta)$ is uniformly continuous on the compact set $\mathcal{Y}$. Thus, there exists a common modulus of continuity $\omega(\,\cdot\,)$ with $\lim_{\eta\searrow 0}\omega(\eta)=0$ such that, for all $\theta\in\Theta$, for all $(a,b)\in\mathcal{A}\times\mathcal{A}$, for all $(\boldsymbol{y},\boldsymbol{z})\in\mathcal{Y}\times\mathcal{Y}$,
$\left\lvert y_{a}c_{b}(\boldsymbol{y},\theta)-z_{a}c_{b}(\boldsymbol{z},\theta)\right\rvert\leq\omega(\max_{b}\lvert y_{b}-z_{b}\rvert).$ (4.6)
Since $\mu$ is BCWE, for all $a,b\in\mathcal{A}$,
$\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}p(\theta)\mu(\boldsymbol{y}\mid\theta)y_{a}c_{a}(\boldsymbol{y},\theta)\leq\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}p(\theta)\mu(\boldsymbol{y}\mid\theta)y_{a}c_{b}(\boldsymbol{y},\theta).$ (4.7)
It follows from Eqs. 4.7, 4.3 and 4.6 that
$\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}p(\theta)\mu(\boldsymbol{y}\mid\theta)\tilde{y}_{a}c_{a}(\tilde{\boldsymbol{y}},\theta)\leq\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}p(\theta)\mu(\boldsymbol{y}\mid\theta)\tilde{y}_{a}c_{b}(\tilde{\boldsymbol{y}},\theta)+2\omega(\eta),$ (4.8)
and thus
$\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}\frac{p(\theta)\mu(\boldsymbol{y}\mid\theta)\tilde{y}_{a}}{\operatorname{\mathsf{Q}}(a)}c_{a}(\tilde{\boldsymbol{y}},\theta)\leq\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}\frac{p(\theta)\mu(\boldsymbol{y}\mid\theta)\tilde{y}_{a}}{\operatorname{\mathsf{Q}}(a)}c_{b}(\tilde{\boldsymbol{y}},\theta)+\frac{2\omega(\eta)}{\operatorname{\mathsf{Q}}(a)}.$ (4.9)
From Eq. 4.3 we have
$\operatorname{\mathsf{Q}}(a)=\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}p(\theta)\mu(\boldsymbol{y}\mid\theta)\tilde{y}_{a}\geq\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}p(\theta)\mu(\boldsymbol{y}\mid\theta)(y_{a}-\eta)\geq\varepsilon_{0}-\eta\geq\frac{\varepsilon_{0}}{2},$ (4.10)
for $\eta\leq\varepsilon_{0}/2$. It follows that
$\frac{2\omega(\eta)}{\operatorname{\mathsf{Q}}(a)}\leq\frac{2\omega(\eta)}{\varepsilon_{0}/2}.$ (4.11)
[…]
Now, take a continuous $\psi\colon\mathcal{Y}\times\Theta\to\mathbb{R}$. Since $\mathcal{Y}\times\Theta$ is compact, $\psi$ is uniformly continuous and admits a modulus of continuity $\omega_{\psi}$ such that for all $\theta\in\Theta$, for all $(a,b)\in\mathcal{A}\times\mathcal{A}$, for all $(\boldsymbol{y},\boldsymbol{z})\in\mathcal{Y}\times\mathcal{Y}$,
$\left\lvert\psi(\boldsymbol{y},\theta)-\psi(\boldsymbol{z},\theta)\right\rvert\leq\omega_{\psi}(\max_{b}\lvert y_{b}-z_{b}\rvert).$ (4.12)
We have for each $\theta\in\Theta$,
$\begin{split}\left\lvert\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}\mu(\boldsymbol{y}\mid\theta)\psi(\tilde{\boldsymbol{y}},\theta)-\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}\mu(\boldsymbol{y}\mid\theta)\psi(\boldsymbol{y},\theta)\right\rvert&\leq\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}\mu(\boldsymbol{y}\mid\theta)\left\lvert\psi(\tilde{\boldsymbol{y}},\theta)-\psi(\boldsymbol{y},\theta)\right\rvert\\ &\leq\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}\mu(\boldsymbol{y}\mid\theta)\omega_{\psi}(\eta)\\ &\leq\omega_{\psi}(\eta),\end{split}$ (4.13)
which is less or equal to $\varepsilon$ for $\eta$ small.
If flows in $\mathcal{Y}^{\ast}$ have rational coefficients with common denominator $K$, then we can choose $\tilde{\boldsymbol{y}}=\boldsymbol{y}$ and $\eta=0$, we have a Bayesian Wardrop 00-equilibrium. Finally, notice that if $\psi$ is $L$-Lipschitz (for the sup norm on flows), then we can choose $\omega_{\psi}(\eta)=L\eta$. Thus, for given $\varepsilon$, the choice of $\eta$ and $\mathcal{I}$ is the same across all such $\psi$. ∎
5 Full Implementation
In this section, we study full implementation of BCWE in games with convex potential.
Definition 5.1.
A basic game $\Gamma=\left(\mathcal{A},\Theta,p,\boldsymbol{c}\right)$ is a potential game if for every $\theta\in\Theta$, there exists an open superset $\widetilde{\mathcal{Y}}$ of $\mathcal{Y}$ in $\mathbb{R}^{\mathcal{A}}$ and a continuously differentiable function $\Phi_{\theta}\colon\widetilde{\mathcal{Y}}\to\mathbb{R}$, called potential, such that, for every $a\in\mathcal{A}$ and $\boldsymbol{y}\in\mathcal{Y}$,
$\frac{\partial\Phi_{\theta}(\boldsymbol{y})}{\partial y_{a}}=c_{a}(\boldsymbol{y},\theta).$ (5.1)
Every game with two actions is a potential game. An important class of potential games is the one of congestion games defined by: a finite set of resources $\mathcal{E}$; a set of actions $\mathcal{A}\subseteq 2^{\mathcal{E}}$; for each resource $e\in\mathcal{E}$, a continuous cost function $c_{e}\colon\mathbb{R}_{+}\times\Theta\to\mathbb{R}_{+}$. The cost of choosing action $a\in\mathcal{A}$ in state $\theta$ is the sum of the costs of all resources used in $a$, $c_{a}(\boldsymbol{y},\theta)\coloneqq\sum_{e\in a}c_{e}(x_{e},\theta)$, where $x_{e}\coloneqq\sum_{a\in\mathcal{A},a\ni e}y_{a}$ is the load on resource $e$ induced by the flow $\boldsymbol{y}$. The potential function is
$\Phi_{\theta}(\boldsymbol{y})=\sum_{e\in\mathcal{E}}\int_{0}^{x_{e}}c_{e}(t,\theta)\ \textup{d}t.$ (5.2)
When for every state, the cost function of each resource is increasing, the potential is convex. If the set of actions is the set of single resources ($\mathcal{A}=\{\{e\},e\in\mathcal{E}\}$) and the cost functions are strictly increasing, the potential is strictly convex.
Proposition 5.2.
Consider a basic game $\Gamma$. If the potential is convex for each state, then for every information structure $\mathcal{I}$, all BWE of $(\Gamma,\mathcal{I})$ have the same cost profile. If the potential is strictly convex for each state, then for every information structure $\mathcal{I}$ there exists a unique BWE outcome of $(\Gamma,\mathcal{I})$.
To prove Proposition 5.2, we first show that a BWE of $(\Gamma,\mathcal{I})$ is equivalent to a Wardrop equilibrium of an auxiliary game with complete information and multiple populations which is also a potential game. If the basic game has a convex potential in each state, then the potential of the auxiliary game is also convex. Uniqueness of Wardrop equilibrium costs of the auxiliary game follows from Koessler et al., (2024, Proposition 4). As a consequence, we get uniqueness of BWE cost profiles for $(\Gamma,\mathcal{I})$. If the basic game has a strictly convex potential in each state, then the potential of the auxiliary game is a strictly convex function of the interim total flow profiles $(\boldsymbol{y}(\boldsymbol{\tau}))_{\boldsymbol{\tau}}$. Uniqueness of the equilibrium interim total flow profile follows then from the fact that the Wardrop equilibrium of the auxiliary multi-population game with complete information are the minimizers of the potential (see Sandholm,, 2001, Proposition 3.1).
Proof of Proposition 5.2.
[…]
$C_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}})\coloneqq\sum_{\theta,\boldsymbol{\tau}^{-k}}p(\theta)\pi\left(\tau^{k},\boldsymbol{\tau}^{-k}\mid\theta\right)c_{a}\left(\left(\sum_{j\in\mathcal{K}}y_{b}^{j}(\tau^{j})\right)_{b\in\mathcal{A}},\theta\right)=\operatorname{\mathsf{P}}(\tau^{k})c_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}}).$ (5.3)
A Wardrop equilibrium of this auxiliary game is identical to a Bayesian Wardrop equilibrium of $(\Gamma,\mathcal{I})$.
Suppose that in each state $\theta$ there is a potential $\Phi_{\theta}$. Then let
$\Phi_{\pi}(\widehat{\boldsymbol{y}})\coloneqq\sum_{\theta,\tau}p(\theta)\pi(\boldsymbol{\tau}\mid\theta)\Phi_{\theta}\left(\left(\sum_{k}y_{a}^{k}(\tau^{k})\right)_{a\in\mathcal{A}}\right)=\sum_{\theta,\tau}p(\theta)\pi(\boldsymbol{\tau}\mid\theta)\Phi_{\theta}(\boldsymbol{y}(\boldsymbol{\tau})),$ (5.4)
where we use the notation $y_{a}(\boldsymbol{\tau})=\sum_{k\in\mathcal{K}}y_{a}^{k}(\tau^{k})$ and $\boldsymbol{y}(\boldsymbol{\tau})=\left(y_{a}(\boldsymbol{\tau})\right)_{a\in\mathcal{A}}.$ The function $\Phi_{\pi}$ is a potential for the auxiliary game: it is immediate to check that
$\frac{\partial\Phi_{\pi}(\widehat{\boldsymbol{y}})}{\partial y_{a}^{k}(\tau^{k})}=C_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}}).$ (5.5)
If $\Phi_{\theta}$ is strictly convex for each $\theta$, then $\Phi_{\pi}$ is a strictly convex function of the vector $(\boldsymbol{y}(\boldsymbol{\tau}))_{\boldsymbol{\tau}}$. By Sandholm, (2001, Proposition 3.1), the equilibrium flows are the minimizers of the potential function. Thus, there is a unique equilibrium vector $(\boldsymbol{y}(\boldsymbol{\tau}))_{\boldsymbol{\tau}}$, and therefore a unique equilibrium outcome, given by (2.7). If $\Phi_{\pi}$ is (weakly) convex, there is a unique equilibrium cost profile $(C_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}}))_{k,\tau^{k},a}$ from Koessler et al., (2024, Proposition 4), and thus a unique BWE cost profile from Eq. 5.3. ∎
The next theorem is a full implementation result for games with strictly convex potential. It shows that if $\mu$ is a BCWE with finite support and rational flows, then there exists a direct information structure such that $\mu$ is implemented as a unique Bayesian Wardrop equilibrium outcome. Further, the information structure is direct and the BWE is obedient. If the support of $\mu$ is not finite or if some flows in the support are not rational, then $\mu$ is close to the unique BWE outcome $\mu^{\ast}$ for some direct information structure $\mathcal{I}$.
Theorem 5.3.
Let $\Gamma$ be a basic game that has a strictly convex potential in each state. Let $\mu$ be a BCWE and let $\psi\colon\mathcal{Y}\times\Theta\to\mathbb{R}$ be a continuous function. For all $\varepsilon>0$, there exists a direct information structure $\mathcal{I}$ such that the Bayesian Wardrop equilibrium outcome $\mu^{\ast}$ of $(\Gamma,\mathcal{I})$ is unique and
$\left\lvert\int\psi(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)-\int\psi(\boldsymbol{y},\theta)\ \textup{d}\mu^{\ast}(\boldsymbol{y}\mid\theta)\right\rvert\leq\varepsilon,\quad\forall\theta\in\Theta.$ (5.6)
In addition, if $\mu$ has finite support with rational flows, then we can choose $\varepsilon=0$ and the obedient outcome $\breve{\mu}_{\mathcal{I}}$ is the unique Bayesian Wardrop equilibrium outcome $\mu^{\ast}$.
As before, $\psi$ represents any possible cost function for the designer. Also, for every $L>0$, we can choose the same $\mathcal{I}$ for all $L$-Lipschitz $\psi$.
The structure of the proof is as follows. Consider the construction given in the proof of Theorem 4.1: We approximate $\mu$ by a distribution over a finite set of rational flows and construct a direct information structure such that the obedient outcome is a Bayesian Wardrop $\alpha$-equilibrium for some $\alpha>0$, and is close to $\mu$. The difficulty is that the obedient outcome need not be an exact Bayesian Wardrop equilibrium and that we don’t know how to construct this BWE explicitly. We show that any Bayesian Wardrop $\alpha$-equilibrium is an $\alpha$-minimizer of the potential of the game with the constructed information structure, and that all $\alpha$-minimizers are $\varepsilon$-close to each other for some $\varepsilon>0$. In particular, the unique BWE outcome $\mu^{\ast}$ is close to the obedient outcome and thus, also close to $\mu$. This reasoning is summarized in the following proposition.
Proposition 5.4.
Let $\Gamma$ be a basic game that has a strictly convex potential in each state. Let $\mu$ be a BCWE and let $\psi\colon\mathcal{Y}\times\Theta\to\mathbb{R}$ be a continuous function. For all $\varepsilon>0$, there exists $\alpha>0$ and a direct information structure $\mathcal{I}$ such that the Bayesian Wardrop equilibrium outcome $\mu^{\ast}$ of $(\Gamma,\mathcal{I})$ is unique, and for all $\tilde{\mu}\in\operatorname{\mathsf{BWE_{\alpha}}}(\mathcal{I})$,
$\left\lvert\int\psi(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)-\int\psi(\boldsymbol{y},\theta)\ \textup{d}\tilde{\mu}(\boldsymbol{y}\mid\theta)\right\rvert\leq\varepsilon,\quad\forall\theta\in\Theta.$ (5.7)
Note that $\mu^{\ast}$ belongs to $\operatorname{\mathsf{BWE_{\alpha}}}(\mathcal{I})$ for all $\alpha\geq 0$.
Proof of Proposition 5.4.
We take up the construction from the proof of Theorem 4.1: fix $\alpha>0$, assume that $\mu$ has finite support, and consider the information structure $\mathcal{I}$ constructed there with parameter $\eta>0$, such that the obedient flow profile $\breve{\boldsymbol{y}}_{\mathcal{I}}$ with outcome $\breve{\mu}_{\mathcal{I}}$ is a Bayesian Wardrop $\alpha$-equilibrium and
$\left\lvert\int\psi(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)-\int\psi(\boldsymbol{y},\theta)\ \textup{d}\breve{\mu}_{\mathcal{I}}(\boldsymbol{y}\mid\theta)\right\rvert\leq\alpha,\quad\forall\theta\in\Theta.$ (5.8)
Claim 5.5.
Suppose that the potential is (weakly) convex in each state. For all $\alpha>0$, any Bayesian Wardrop $\alpha$-equilibrium of the game with information structure $\mathcal{I}$, is an $\alpha$-minimizer of the associated potential.
Proof of 5.5.
Consider an arbitrary Bayesian Wardrop $\alpha$-equilibrium $\widehat{\boldsymbol{y}}$. For all $k\in\mathcal{K}$, for all $\tau^{k}\in\mathcal{T}^{k}$, and for all $a,b\in\mathcal{A}$, if $y_{a}^{k}(\tau^{k})>0$ then,
$C_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}})\leq C_{b}^{k,\tau^{k}}(\widehat{\boldsymbol{y}})+\alpha\operatorname{\mathsf{P}}(\tau^{k})\leq C_{b}^{k,\tau^{k}}(\widehat{\boldsymbol{y}})+\alpha,$ (5.9)
where $C_{a}^{k,\tau^{k}}$ is defined as in (5.3). This implies that, for all $k\in\mathcal{K}$, for all $\tau^{k}\in\mathcal{T}^{k}$,
$\sum_{a\in\mathcal{A}}\frac{y_{a}^{k}(\tau^{k})}{\gamma^{k}}C_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}})\leq\min_{b\in\mathcal{A}}C_{b}^{k,\tau^{k}}(\widehat{\boldsymbol{y}})+\alpha=\min_{\widehat{\boldsymbol{z}}}\sum_{a\in\mathcal{A}}\frac{z_{a}^{k}(\tau^{k})}{\gamma^{k}}C_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}})+\alpha.$ (5.10)
That is, for all $k\in\mathcal{K}$, for all $\tau^{k}\in\mathcal{T}^{k}$ and all $\widehat{\boldsymbol{z}}$,
$\sum_{a\in\mathcal{A}}\left(y_{a}^{k}(\tau^{k})-z_{a}^{k}(\tau^{k})\right)C_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}})\leq\gamma^{k}\alpha.$ (5.11)
Summing up over $k$ and $\tau^{k}$, we get,
$\left\langle\nabla\Phi_{\pi}(\widehat{\boldsymbol{y}}),\widehat{\boldsymbol{y}}-\widehat{\boldsymbol{z}}\right\rangle=\sum_{k}\sum_{\tau^{k}}\sum_{a\in\mathcal{A}}\left(y_{a}^{k}(\tau^{k})-z_{a}^{k}(\tau^{k})\right)C_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}})\leq\sum_{k}\gamma^{k}\alpha=\alpha,$ (5.12)
and since $\Phi_{\pi}$ is convex, this gives
$\Phi_{\pi}(\widehat{\boldsymbol{y}})-\Phi_{\pi}(\widehat{\boldsymbol{z}})\leq\left\langle\nabla\Phi_{\pi}(\widehat{\boldsymbol{y}}),\widehat{\boldsymbol{y}}-\widehat{\boldsymbol{z}}\right\rangle\leq\alpha.$ (5.13)
Thus $\Phi_{\pi}(\widehat{\boldsymbol{y}})\leq\min\Phi_{\pi}+\alpha$. ∎
From strict convexity, let $(\boldsymbol{y}^{*}(\boldsymbol{\tau}))_{\boldsymbol{\tau}}$ be the unique Bayesian Wardrop equilibrium flow vector. We argue now that for any $\alpha$-minimizer $\widehat{\boldsymbol{y}}$ of $\Phi_{\pi}$, the flow vector $(\boldsymbol{y}(\boldsymbol{\tau}))_{\boldsymbol{\tau}}$ is close to $(\boldsymbol{y}^{*}(\boldsymbol{\tau}))_{\boldsymbol{\tau}}$.
Claim 5.6.
For all $\delta>0$, there exists $\alpha>0$ such that for all $\widehat{\boldsymbol{y}}$, $\Phi_{\pi}(\widehat{\boldsymbol{y}})\leq\min\Phi_{\pi}+\alpha$ implies that for all $a$ and all $\boldsymbol{\tau}$, $\left\lvert y_{a}(\boldsymbol{\tau})-y^{*}_{a}(\boldsymbol{\tau})\right\rvert\leq\delta$.
Proof of 5.6.
By contradiction, suppose that there exists $\delta>0$ such that for all $\alpha=1/n$ with $n\in\mathbb{N}$, there exists $\widehat{\boldsymbol{y}}_{n}$ with $\Phi_{\pi}(\widehat{\boldsymbol{y}}_{n})\leq\min\Phi_{\pi}+1/n$ and $\max_{a,\boldsymbol{\tau}}\left\lvert y_{n,a}(\boldsymbol{\tau})-y^{*}_{a}(\boldsymbol{\tau})\right\rvert>\delta$. By compactness, take a subsequence and assume that $\widehat{\boldsymbol{y}}_{n}\to\widehat{\boldsymbol{y}}_{\infty}$ as $n\to\infty$. Letting $n\to\infty$, we get that $\Phi_{\pi}(\widehat{\boldsymbol{y}}_{\infty})\leq\min\Phi_{\pi}$ and $\max_{a,\boldsymbol{\tau}}\left\lvert\widehat{y}_{\infty,a}(\boldsymbol{\tau})-y^{*}_{a}(\boldsymbol{\tau})\right\rvert>\delta$. This contradicts the uniqueness of $y^{*}_{a}(\boldsymbol{\tau})$. ∎
Take a Bayesian Wardrop $\alpha$-equilibrium with flow vector $(\tilde{\boldsymbol{y}}(\boldsymbol{\tau}))_{\boldsymbol{\tau}}$ and outcome $\tilde{\mu}$. We have for each $\theta\in\Theta$,
$\begin{split}\left\lvert\int\psi(\boldsymbol{y},\theta)\ \textup{d}\mu^{\ast}(\boldsymbol{y}\mid\theta)-\int\psi(\boldsymbol{y},\theta)\ \textup{d}\tilde{\mu}(\boldsymbol{y}\mid\theta)\right\rvert&=\left\lvert\sum_{\boldsymbol{\tau}}\pi(\boldsymbol{\tau}\mid\theta)\psi(\boldsymbol{y}^{*}(\boldsymbol{\tau}),\theta)-\sum_{\boldsymbol{\tau}}\pi(\boldsymbol{\tau}\mid\theta)\psi(\tilde{\boldsymbol{y}}(\boldsymbol{\tau}),\theta)\right\rvert\\ &\leq\sum_{\boldsymbol{\tau}}\pi(\boldsymbol{\tau}\mid\theta)\left\lvert\psi(\boldsymbol{y}^{*}(\boldsymbol{\tau}),\theta)-\psi(\tilde{\boldsymbol{y}}(\boldsymbol{\tau}),\theta)\right\rvert\\ &\leq\omega_{\psi}(\delta).\end{split}$
Summing up, for all $\theta\in\Theta$, we have the following inequalities. By construction, we have
$\left\lvert\int\psi(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)-\int\psi(\boldsymbol{y},\theta)\ \textup{d}\breve{\mu}_{\mathcal{I}}(\boldsymbol{y}\mid\theta)\right\rvert\leq\alpha.$ (5.14)
Moreover, because the obedient interim profile is a Bayesian Wardrop $\alpha$-equilibrium, we have
$\left\lvert\int\psi(\boldsymbol{y},\theta)\ \textup{d}\breve{\mu}_{\mathcal{I}}(\boldsymbol{y}\mid\theta)-\int\psi(\boldsymbol{y},\theta)\ \textup{d}\mu^{\ast}(\boldsymbol{y}\mid\theta)\right\rvert\leq\omega_{\psi}(\delta)$ (5.15)
and for any Bayesian Wardrop $\alpha$-equilibrium:
$\left\lvert\int\psi(\boldsymbol{y},\theta)\ \textup{d}\mu^{\ast}(\boldsymbol{y}\mid\theta)-\int\psi(\boldsymbol{y},\theta)\ \textup{d}\tilde{\mu}(\boldsymbol{y}\mid\theta)\right\rvert\leq\omega_{\psi}(\delta).$ (5.16)
It follows that
$\left\lvert\int\psi(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)-\int\psi(\boldsymbol{y},\theta)\ \textup{d}\tilde{\mu}(\boldsymbol{y}\mid\theta)\right\rvert\leq\alpha+2\omega_{\psi}(\delta),$ (5.17)
which is less or equal to $\varepsilon$ for $\alpha,\delta$ small enough. ∎
Proof of Theorem 5.3.
The result follows from the fact that the unique equilibrium outcome $\mu^{\ast}$ is an Bayesian Wardrop $\alpha$-equilibrium outcome for any $\alpha>0$. As in the proof of Theorem 4.1, if $\mu$ has finite support with rational flows, then we can choose $\varepsilon=0$, and the obedient outcome $\breve{\mu}_{\mathcal{I}}$ is the unique Bayesian Wardrop equilibrium outcome $\mu^{\ast}$. ∎
The next theorem shows that in games with weakly convex potential, we can fully implement the ex ante expected social cost of any Bayes correlated Wardrop equilibrium, where the social cost in state $\theta$ given flows $\boldsymbol{y}$ is:
$\operatorname{\mathsf{SC}}(\boldsymbol{y},\theta)=\sum_{a}y_{a}c_{a}(\boldsymbol{y},\theta).$ (5.18)
Theorem 5.7.
Let $\Gamma$ be a basic game that has a convex potential in each state and let $\mu$ be a BCWE. For all $\varepsilon>0$, there exists a direct information structure $\mathcal{I}$ such that for all Bayesian Wardrop equilibrium outcome $\mu^{\ast}$ of $(\Gamma,\mathcal{I})$,
$\left\lvert\sum_{\theta}p(\theta)\int\operatorname{\mathsf{SC}}(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)-\sum_{\theta}p(\theta)\int\operatorname{\mathsf{SC}}(\boldsymbol{y},\theta)\ \textup{d}\mu^{\ast}(\boldsymbol{y}\mid\theta)\right\rvert\leq\varepsilon.$ (5.19)
As before, if $\mu$ has finite support and rational flows, we can choose $\varepsilon=0$ and we know that the obedient outcome $\breve{\mu}_{\mathcal{I}}$ is a Bayesian Wardrop equilibrium outcome.
Theorem 5.7 applies to all congestion games with increasing costs. For instance, in Example 1, it implies that there exists an information structure (the one described on page 1) that implements a social cost-minimizing BCWE in every Bayesian Wardrop equilibrium of the basic game extended with this information structure. However, Theorem 5.7 does not apply to Example 2, which is a potential game but lacks convexity. In this case, Theorem 4.1 implies that there exists an information structure (the one described on page 2) that implements the social cost-minimizing BCWE in a Bayesian Wardrop equilibrium of the basic game extended with this information structure, but other, less efficient equilibria also exist. [^7] Theorem 5.7 also does not apply to games with positive payoff externalities (i.e., strictly decreasing costs), such as those in Li et al., (2023). However, Li et al., (2023) provide an example showing that full implementation may still be possible with complex information structures involving an infinite set of signals.
Proof of Theorem 5.7.
Assume that $\mu$ is a BCWE with finite support, choose the function $\psi$ as the social cost function $\operatorname{\mathsf{SC}}$, and fix $\alpha>0$. Consider the information structure $\mathcal{I}$ constructed in the proof of Theorem 4.1 with parameter $\eta>0$, such that the obedient flow profile $\breve{\boldsymbol{y}}_{\mathcal{I}}$ with outcome $\breve{\mu}_{\mathcal{I}}$ is a Bayesian Wardrop $\alpha$-equilibrium and
$\left\lvert\int\operatorname{\mathsf{SC}}(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)-\int\operatorname{\mathsf{SC}}(\boldsymbol{y},\theta)\ \textup{d}\breve{\mu}_{\mathcal{I}}(\boldsymbol{y}\mid\theta)\right\rvert\leq\alpha,\quad\forall\theta\in\Theta.$ (5.20)
We know that the set of Wardrop equilibria of the associated multi-population game with complete information is the set of minimizers of the potential (Sandholm,, 2001, Proposition 3.1). Since there is a convex potential in each state, the potential defined in Eq. 5.4 is a convex function and the set of Wardrop equilibria is convex. Also, from Koessler et al., (2024, Proposition 4), all Wardrop equilibria have the same cost profile. That is, for any Wardrop equilibrium flow vectors $\boldsymbol{y}^{*}=(\boldsymbol{y}^{*}(\boldsymbol{\tau}))_{\boldsymbol{\tau}}$ and $\boldsymbol{y}^{**}=(\boldsymbol{y}^{**}(\boldsymbol{\tau}))_{\boldsymbol{\tau}}$, for every $k,\tau^{k}$,
$C_{a^{*}}^{k,\tau^{k}}(\boldsymbol{y}^{*})=C_{a^{**}}^{k,\tau^{k}}(\boldsymbol{y}^{**}),$ (5.21)
where $a^{*}$ (resp. $a^{**}$) is any action in the support of $\boldsymbol{y}^{*}$ (resp. $\boldsymbol{y}^{**}$) for population $(k,\tau^{k})$. It follows that all Wardrop equilibria have the same total cost, where the total cost of an interim flow profile $\widehat{\boldsymbol{y}}$ is,
$\operatorname{\mathsf{TC}}(\widehat{\boldsymbol{y}})\coloneqq\sum_{k,\tau^{k}}\sum_{a}y_{a}^{k}(\tau^{k})C_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}}).$ (5.22)
Claim 5.8.
For any interim flow profile $\widehat{\boldsymbol{y}}$, the total cost is the expectation of the social cost function.
Proof.
From (5.3) we have that the total cost in the population game is
$\begin{split}\operatorname{\mathsf{TC}}(\widehat{\boldsymbol{y}})&=\sum_{k,\tau^{k}}\sum_{a}y_{a}^{k}(\tau^{k})\sum_{\theta,\boldsymbol{\tau}^{-k}}p(\theta)\pi\left(\tau^{k},\boldsymbol{\tau}^{-k}\mid\theta\right)c_{a}(\boldsymbol{y}(\boldsymbol{\tau}),\theta)\\ &=\sum_{\theta}p(\theta)\sum_{\boldsymbol{\tau}}\pi\left(\boldsymbol{\tau}\mid\theta\right)\sum_{a}\sum_{k}y_{a}^{k}(\tau^{k})c_{a}(\boldsymbol{y}(\boldsymbol{\tau}),\theta)\\ &=\sum_{\theta}p(\theta)\sum_{\boldsymbol{\tau}}\pi\left(\boldsymbol{\tau}\mid\theta\right)\sum_{a}y_{a}(\boldsymbol{\tau})c_{a}(\boldsymbol{y}(\boldsymbol{\tau}),\theta)\\ &=\sum_{\theta}p(\theta)\sum_{\boldsymbol{\tau}}\pi\left(\boldsymbol{\tau}\mid\theta\right)\operatorname{\mathsf{SC}}(\boldsymbol{y},\theta).\end{split}$ (5.23)
Corollary 5.9.
The expected social cost is the same in all Bayesian Wardrop equilibrium outcomes of $\Gamma$ with information structure $\mathcal{I}$.
From 5.5, any Bayesian Wardrop $\alpha$-equilibrium of the game with information structure $\mathcal{I}$, is a $\alpha$-minimizer of the associated potential. To complete the proof, we use the following argument:
Claim 5.10.
For all $\delta>0$, there exists $\alpha>0$ such that for all $\widehat{\boldsymbol{y}}$, $\Phi_{\pi}(\widehat{\boldsymbol{y}})\leq\min\Phi_{\pi}+\alpha$ implies
$\left\lvert\operatorname{\mathsf{TC}}(\widehat{\boldsymbol{y}})-\operatorname{\mathsf{TC}}(\widehat{\boldsymbol{y}}^{*})\right\rvert\leq\delta,$ (5.24)
where $\widehat{\boldsymbol{y}}^{*}$ is any Bayesian Wardrop equilibrium of $\Gamma$ with information structure $\mathcal{I}$.
Proof of 5.10.
By contradiction, suppose that there exists $\delta>0$ such that for all $\alpha=1/n$ with $n\in\mathbb{N}$, there exists $\widehat{\boldsymbol{y}}_{n}$ with $\Phi_{\pi}(\widehat{\boldsymbol{y}}_{n})\leq\min\Phi_{\pi}+1/n$ and $\left\lvert\operatorname{\mathsf{TC}}(\widehat{\boldsymbol{y}}_{n})-\operatorname{\mathsf{TC}}(\widehat{\boldsymbol{y}}^{*})\right\rvert>\delta$. By compactness, take a subsequence and assume that $\widehat{\boldsymbol{y}}_{n}\to\widehat{\boldsymbol{y}}_{\infty}$ as $n\to\infty$. Letting $n\to\infty$, we get that $\Phi_{\pi}(\widehat{\boldsymbol{y}}_{\infty})\leq\min\Phi_{\pi}$ and $\left\lvert\operatorname{\mathsf{TC}}(\widehat{\boldsymbol{y}}_{\infty})-\operatorname{\mathsf{TC}}(\widehat{\boldsymbol{y}}^{*})\right\rvert>\delta$. This is a contradiction since $\widehat{\boldsymbol{y}}_{\infty}$ would then be a minimizer of the potential, thus a Bayesian Wardrop equilibrium, with a different total cost. ∎
This concludes the proof of Theorem 5.7. ∎
6 List of Symbols
$\mathcal{A}$ set of action profiles, defined in Definition 2.1
$a$ action
$\boldsymbol{a}$ action profile
$b$ action
$\operatorname{\mathsf{BWE_{\varepsilon}}}(\mathcal{I})$ set of Bayesian Wardrop $\varepsilon$-equilibrium outcomes of $\Gamma$ with information structure $\mathcal{I}$
$c_{a}$ cost of action $a$, defined in Definition 2.1
$c_{a}^{k,\tau^{k}}$ cost functions of population $(k,\tau^{k})$, defined in (5.3)
$\boldsymbol{c}$ $(c_{a})_{a\in\mathcal{A}}$, defined in Definition 2.1
$\mathcal{I}$ $\mathcal{I}=\left(\boldsymbol{\gamma},\mathcal{T},\pi\right)$ , information structure, defined in Definition 2.3
$k$ population
$\mathcal{K}$ population set
$p$ prior distribution on $\Theta$, defined in Definition 2.1
$\operatorname{\mathsf{P}}(\tau^{k})$ $\sum_{\theta\in\Theta}\sum_{\boldsymbol{\tau}^{-k}\in\mathcal{T}^{-k}}p(\theta)\pi(\tau^{k},\boldsymbol{\tau}^{-k}\mid\theta)$, total probability of type $\tau^{k}$
$\operatorname{\mathsf{Q}}^{k}(a)$ $\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}p(\theta)\mu(\boldsymbol{y}\mid\theta)y_{a}^{k}$, defined in (4.5)
$\mathcal{T}$ $\mathcal{T}=\times_{k\in\mathcal{K}}\mathcal{T}^{k}$, defined in Definition 2.3
$\mathcal{T}^{k}$ set of types for population $k$, defined in Definition 2.3
$\mathcal{T}^{-k}$$\times_{j\neq k}\mathcal{T}^{j}$
$\operatorname{\mathsf{WE}}(\Gamma)$ Wardrop equilibria of $\Gamma$
$y_{a}^{k}(\tau^{k})$ flow on action $a$ of population $k$, when type $\tau^{k}$ is observed
$y_{a}(\boldsymbol{\tau})$ $\sum_{k\in\mathcal{K}}y_{a}^{k}(\tau^{k})$, interim total flow on action $a$
$\breve{\boldsymbol{y}}_{\mathcal{I},a}^{k}(\tau^{k})$ $\gamma^{k}\cdot\operatorname{\mathds{1}}\left\{\tau^{k}=a\right\}$
$\boldsymbol{y}$ flow vector
$\boldsymbol{y}^{\ast}$ equilibrium flow vector
$\boldsymbol{y}^{k}(\tau^{k})$ $\left(y_{a}^{k}(\tau^{k})\right)_{a\in\mathcal{A}}\in\Delta_{\gamma^{k}}(\mathcal{A})$, interim flow profile of population $k$ given type $\tau^{k}$
$\boldsymbol{y}(\boldsymbol{\tau})$ $\left(y_{a}(\boldsymbol{\tau})\right)_{a\in\mathcal{A}}\in\Delta(\mathcal{A})$, interim total flow profile given $\boldsymbol{\tau}$
$\widehat{\boldsymbol{y}}$ $\left(\boldsymbol{y}^{k}(\tau^{k})\right)_{k,\tau^{k}}\in\bigtimes_{k\in\mathcal{K}}\left[\bigtimes_{\tau^{k}\in\mathcal{T}^{k}}\Delta_{\gamma^{k}}(\mathcal{A})\right]$, interim flow profile, defined in (2.3)
$\breve{\boldsymbol{y}}_{\mathcal{I}}$ obedient interim flow profile
$\mathcal{Y}$ $\Delta_{\gamma}(\mathcal{A})$, set of flow profiles, defined in Definition 2.1
$\gamma$ $\sum_{k\in\mathcal{K}}\gamma^{k}$ total population size, defined in Definition 2.1
$\gamma^{k}$ size of population $k$, defined in Definition 2.3
$\boldsymbol{\gamma}$ vector of population sizes, defined in Definition 2.3
$\Gamma$ $\left(\gamma,\mathcal{A},\Theta,p,\boldsymbol{c}\right)$, Bayesian nonatomic game, defined in Definition 2.1
$\delta_{a}$ Dirac measure on $a$
$\Delta(A)$ $\Delta_{1}(A)$, simplex of probability measures on $A$
$\Delta_{\gamma}(A)$$\left\{\boldsymbol{y}\in\mathbb{R}^{A}\colon\forall a\in A,y_{a}\geq 0,\sum_{a\in A}y_{a}=\gamma\right\}$, defined in (1.1)
$\theta$ state
$\Theta$ state space, defined in Definition 2.1
$\mu$ outcome of the game, defined in Definition 2.1
$\mu^{\ast}$ Bayesian Wardrop equilibrium outcome
$\breve{\mu}_{\mathcal{I}}$ outcome associated to the obedient interim flow profile
$\pi$ mapping from $\Theta$ to $\Delta(\mathcal{T})$, defined in Definition 2.3
$\tau^{k}$ possible type for population $k$
$\boldsymbol{\tau}=(\tau^{k})_{k\in\mathcal{K}}$ profile of types
$\Phi_{\theta}$ potential, defined in Definition 5.1
$\omega(\,\cdot\,)$ modulus of continuity
$\operatorname{\mathds{1}}$ indicator function