### v2

Abstract
This paper studies the implementation of Bayes correlated equilibria in symmetric Bayesian games with nonatomic players, using direct information structures and obedient strategies. The main results demonstrate full implementation in a class of games with negative payoff externalities, such as congestion and Cournot games. Specifically, if the game admits a strictly concave potential in every state, then for every Bayes correlated equilibrium outcome with finite support and rational action distributions, there exists a direct information structure that implements this outcome under all equilibria. When the potential is weakly concave, we show that all equilibria implement the same expected total payoff. Additionally, all Bayes correlated equilibria, including those with infinite support or irrational action distributions, are approximately implemented.
Keywords: Bayes correlated equilibrium, congestion games, aggregative games, information design, nonatomic games, population games, potential games.
1 Introduction
The fundamental equilibrium concepts used in information design for incomplete-information games with finitely many players, are Bayes Nash equilibrium and Bayes correlated equilibrium. The use of these concepts in games with a large number of players is often cumbersome and computationally intractable. Yet, large games with symmetries can be approximated by nonatomic games, whose natural equilibrium concept in the case of complete information—Wardrop equilibrium—is based on distributions over actions (called flows), rather than on the behavior of single players. This technique was recently extended to incomplete-information nonatomic games, and the concepts of Bayesian Wardrop equilibrium (BWE) and Bayes correlated Wardrop equilibrium (BCWE) were introduced and analyzed (see Koessler et al.,, 2025).
In this paper, we study partial and full implementation through information design in symmetric nonatomic games with incomplete information. In these, there is a payoff-relevant state, to which information will be correlated by a perfectly informed designer whose goal is—for instance—to maximize total welfare. In this context, the concept of BCWE characterizes the set of state-dependent probability distributions over flows that can be induced in an incentive-compatible way by some information structure. Furthermore, regardless of the designer’s objective, a designer-optimal BCWE can be formulated as a solution to an optimization program with a bounded number of variables and constraints, determined by the number of actions and states.
We show that approaching information design through the lens of nonatomic games provides important insights, particularly for achieving full implementation through direct recommendation systems. There is full implementation when all equilibria implement the same outcome for the chosen information structure. This property is crucial because it ensures robustness of the information design problem to adversarial equilibrium selection. When only some equilibria induce the desired outcome, we talk about partial implementation.
Our basic game is a nonatomic game with symmetric and incomplete information. Our analysis extends the class of population games studied by Sandholm, (2001, 2010) to the Bayesian setting. [^4] Sandholm, (2001, 2010) allowed for multiple populations with different preferences and sets of feasible actions. For notational simplicity, in this paper we assume that the basic game involves only one population, so players are ex-ante symmetric. At the cost of cumbersome notation, our results can be generalized to multiple populations. We extend this basic game to asymmetric information by considering a family of simple information structures (as defined in Wu et al.,, 2021): The set of players is partitioned into a finite number of populations; within any given population, all players receive the same signal. Such an information structure is a direct information structure if the set of possible signals is identical to the set of possible actions. A BWE of the game extended with an information structure is said to be obedient if all players in all populations follow their recommended action.
We first show that every BCWE with finite support and rational flows can be partially implemented as an obedient BWE with an appropriate direct information structure (Theorem˜4.1). Hence, a version of the revelation principle (Myerson,, 1982) applies to symmetric nonatomic Bayesian games. Next, we turn to full implementation, which ensures that every BWE outcome of the extended game leads to the desired BCWE. We show that if the basic game admits a strictly concave potential in each state, then the BWE outcome is unique for every information structure (Theorem˜5.5). Consequently, information design becomes very effective in nonatomic games with a strictly concave potential because every BCWE with finite support and rational flows can be fully implemented, i.e., there exists a direct information structure such that the BCWE is implemented, regardless of the BWE selection.
If in each state the basic game admits a potential that is just weakly concave, full implementation of a given outcome may not be achieved, yet we get full payoff implementation. We show that for every BCWE with finite support and rational flows, there exists a direct information structure such that expected total payoff is the same as the BCWE expected total payoff regardless of the BWE (Theorem˜5.9). In particular, the expected payoff of the total payoff-maximizing BCWE can be fully implemented.
[…]
Our full (payoff) implementation results apply, in particular, to congestion games with decreasing payoff functions and to Cournot games. A congestion game has a strictly concave potential if the set of actions is equal to the set of resources and the payoff of using a resource is strictly decreasing in the number of users. When actions are general feasible subsets of resources and payoffs are added across resources, congestion games admit a concave potential, provided that the payoff of using each resource is weakly decreasing in the number of users. A Cournot game admits a concave potential if inverse demand is strictly decreasing, and the potential is strictly concave when possible quantities for each firm are binary (e.g., entry and no entry).
1.1 Related Literature
The notion of Wardrop equilibrium originated from the work of Wardrop, (1952) on traffic flow theory, where he formulated principles describing user-optimal and system-optimal routing in transportation networks. Beckmann et al., (1956) proved that a Wardrop equilibrium is the solution of a convex optimization problem. This idea was framed in the language of potential games (see, among others, Sandholm,, 2001, Roughgarden,, 2007, Roughgarden and Tardos,, 2007), generalizing the case of a finite number of players (Rosenthal,, 1973, Monderer and Shapley,, 1996) to the nonatomic case. A general framework for nonatomic potential games, extending beyond routing games, was studied by Sandholm, (2001), Hofbauer and Sandholm, (2009), Sandholm, (2010). The class of aggregative games—subclass of potential games—was implicitly introduced by Dubey et al., (1980) and formalized by Corchón, (1994). Lahkar, (2017) studied this class in games with nonatomic players and a finite set of actions.
Koessler et al., (2025) extended this framework to nonatomic Bayesian games by studying the notions of correlated and Bayes correlated Wardrop equilibria. They showed that Bayes correlated Wardrop equilibria emerge as limits of action flows induced by Bayes correlated equilibria (Bergemann and Morris,, 2016) in finite-player games. [^5] In Koessler et al., (2025) players minimize cost functions, whereas here players maximize payoff functions, which are simply the negatives of cost functions. Related nonatomic versions of correlated and Bayes correlated equilibria were studied, e.g., by Díaz et al., (2009), Tavafoghi and Teneketzis, (2018), and Zhu and Savla, (2022). Acemoglu et al., (2018) showed the relevance of information design in routing games, and proved that some classical paradoxes in routing games can be expressed in terms of negative value of information. Among the recent contributions on the topic, the reader is referred, for instance, to Massicot and Langbort, (2022), Cianfanelli et al., (2023), Ambrogio et al., (2024), and Griesbach et al., (2024). In a concurrent work focusing on congestion games and using distributional approach, Massicot, (2025, chapter 5) considers infinite information structures and obtains implementation results related to ours.
To the best of our knowledge, our full implementation results with direct information structures have no analog in the literature on information design, where—even for binary action games—full implementation typically relies on complex indirect information structures, often with infinite type spaces. Existing tools and recent results in this area include, for example, Mathevet et al., (2020), Dworczak and Pavan, (2022), Hoshino, (2022), Inostroza and Pavan, (2025), Li et al., (2023), and Morris et al., (2024). The literature on information design is surveyed in Bergemann and Morris, (2019), Kamenica, (2019) and Forges, (2020). Dughmi, (2017) surveys the algorithmic literature on information design.
1.2 Organization of the Paper
Section˜2 presents the model, including the basic game, its Bayes correlated Wardrop equilibria, information structures with finite populations and public signals within populations, and Bayesian Wardrop ($\varepsilon$-)equilibria of the basic game extended with an information structure. Section˜3 provides two simple examples and illustrates how to implement welfare-maximizing BCWE. Our partial and full implementation results are presented in Sections˜4 and 5, respectively. In the Appendix, Section˜6 identifies a class of games for which deterministic flows maximize the designer’s payoff function. Appendix˜A contains a list of symbols and acronyms used throughout the paper.
1.3 Notation
Given any finite set $A$ and $\gamma>0$, we define
$\Delta_{\gamma}(A)\coloneqq\left\{\boldsymbol{y}=\left(y_{a}\right)_{a}\colon y_{a}\geq 0,\sum_{a\in A}y_{a}=\gamma\right\}.$ (1.1)
For the sake of simplicity, we use the shorthand $\Delta(A)$ to indicate $\Delta_{1}(A)$, which is just the simplex of probability measures on $A$. For a compact set $\mathcal{Y}$, $\Delta{(\mathcal{Y})}$ denotes the set of Borel probability measures on $\mathcal{Y}$. For $h\in\mathbb{N}$, the symbol $C^{h}$ denotes the class of functions having an $h$-th derivative that is continuous in its domain.
2 The Model
We consider a basic game $\Gamma$, which is a symmetric nonatomic Bayesian game with a finite set of actions, and a finite set of states. Without any loss of generality, throughout the paper, the total mass of players is assumed to be $1$. To define the basic game, we need the concept of flow profile, which is the way the mass of players is distributed over the possible actions. Given a finite set of actions $\mathcal{A}$, a flow profile $\boldsymbol{y}\coloneqq(y_{a})_{a\in\mathcal{A}}\in\mathcal{Y}$, is a nonnegative vector in $\mathcal{Y}\coloneqq\Delta(\mathcal{A})$, and $y_{a}$ is the flow on action $a$.
Definition 2.1.
A basic game $\Gamma=\left(\mathcal{A},\Theta,p,\boldsymbol{u}\right)$ is given by a finite set of actions $\mathcal{A}$, a finite set of states $\Theta$, a full-support probability distribution $p\in\Delta(\Theta)$ over states, and a profile of payoff functions $\boldsymbol{u}=(u_{a})_{a\in\mathcal{A}}$, where, for each $a\in\mathcal{A}$, $u_{a}\colon\mathcal{Y}\times\Theta\to\mathbb{R}$ is continuous.
For every $a\in\mathcal{A}$, $\theta\in\Theta$ and $\boldsymbol{y}\in\mathcal{Y}$, $u_{a}(\boldsymbol{y},\theta)$ is the payoff of action $a$ when the state is $\theta$ and the flow profile in that state is $\boldsymbol{y}$. A transition probability $\mu\colon\Theta\to\Delta(\mathcal{Y})$ that associates a distribution over flows to each state is called an outcome of the game.
The notion of Bayes correlated Wardrop equilibrium in the next definition is the analog of the notion of Bayes correlated equilibrium (Bergemann and Morris,, 2016) in the context of nonatomic games (see Koessler et al.,, 2025).
Definition 2.2.
A Bayes correlated Wardrop equilibrium (BCWE) of $\Gamma$ is an outcome $\mu:\Theta\to\Delta(\mathcal{Y})$ such that
$\sum_{\theta\in\Theta}p(\theta)\int y_{a}u_{a}(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)\geq\sum_{\theta\in\Theta}p(\theta)\int y_{a}u_{b}(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta),\quad\forall\,a,b\in\mathcal{A}.$ (2.1)
For each state $\theta\in\Theta$, a flow profile $\boldsymbol{y}$ is drawn at random according to the distribution $\mu(\,\cdot\mid\theta)$. If we divide the two sides of (2.1) by the ex-ante expected mass of players choosing $a$, then the left-hand-side is proportional to the expected payoff of playing $a$, conditionally of being recommended $a$, and the right-hand side is proportional to the expected payoff of playing $b$, conditionally of being recommended $a$. The corresponding outcome is a BCWE if, for every action $a$ that has strictly positive flow in at least one state, conditionally on being assigned $a$, the expected payoff of playing $a$ is greater or equal to the expected payoff of playing any other action $b$.
For example, suppose that for every $\theta$, $\boldsymbol{y}(\theta)$ is a Wardrop equilibrium of the complete information game $(\mathcal{A},(u_{a}(\,\cdot\,,\theta))_{a\in\mathcal{A}})$, i.e.,
$\forall a,b\in\mathcal{A},\ y_{a}(\theta)>0\implies u_{a}(\boldsymbol{y}(\theta),\theta)\geq u_{b}(\boldsymbol{y}(\theta),\theta).$ (2.2)
Then $\mu(\theta)=\delta_{\boldsymbol{y}(\theta)}$ is a “fully revealing” BCWE.
As another example, if $\bar{\boldsymbol{y}}$ is a Wardrop equilibrium of the average game with complete information $(\mathcal{A},(\bar{u}_{a})_{a\in\mathcal{A}})$, with $\bar{u}_{a}(\,\cdot\,)=\sum_{\theta}p(\theta)u_{a}(\,\cdot\,,\theta)$, then $\mu(\theta)=\delta_{\bar{\boldsymbol{y}}}$ is a “non-revealing” BCWE. This corresponds to an equilibrium outcome of the basic game with no information.
We now consider finite information structures and Bayesian Wardrop equilibria as defined in Wu et al., (2021). Using finite information structure has the following merits: On one hand this avoids measurability issues related to private signals for uncountable sets of players. On the other hand, the approach is flexible enough to guarantee the possibility of arbitrarily fine information diversification.
Definition 2.3.
An information structure $\mathcal{I}=\left(\boldsymbol{\gamma},\mathcal{T},\pi\right)$ for $\Gamma$ is given by:
  • a vector $\boldsymbol{\gamma}=(\gamma^{k})_{k\in\mathcal{K}}\in\Delta(\mathcal{K})$, where $\mathcal{K}$ is a finite set of populations and, for every $k\in\mathcal{K}$, $\gamma^{k}>0$ is the size of population $k$;
  • for each population $k$, a finite set of types $\mathcal{T}^{k}$;
  • a mapping $\pi\colon\Theta\to\Delta(\mathcal{T})$, where $\mathcal{T}=\times_{k\in\mathcal{K}}\mathcal{T}^{k}$.
The game $\Gamma$ extended by information structure $\mathcal{I}$, denoted by $(\Gamma,\mathcal{I})$, can be described as follows: for each state $\theta\in\Theta$, a type profile $\boldsymbol{\tau}\in\mathcal{T}$ is drawn with probability $\pi(\boldsymbol{\tau}\mid\theta)$ and, for every $k$, the entire population $k$ observes the type $\tau^{k}\in\mathcal{T}^{k}$. Then, actions are chosen.
For all $k\in\mathcal{K}$, $\tau^{k}\in\mathcal{T}^{k}$ and $a\in\mathcal{A}$, we let $y_{a}^{k}(\tau^{k})$ denote the flow on action $a$ of population $k$, when type $\tau^{k}$ is observed. The interim flow profile of population $k$ given type $\tau^{k}$ is $\boldsymbol{y}^{k}(\tau^{k})=\left(y_{a}^{k}(\tau^{k})\right)_{a\in\mathcal{A}}\in\Delta_{\gamma^{k}}(\mathcal{A})$. The list
$\widehat{\boldsymbol{y}}=\left(\boldsymbol{y}^{k}(\tau^{k})\right)_{k,\tau^{k}}\in\bigtimes_{k\in\mathcal{K}}\left[\bigtimes_{\tau^{k}\in\mathcal{T}^{k}}\Delta_{\gamma^{k}}(\mathcal{A})\right],$ (2.3)
of interim flows profiles of all populations is called interim flow profile. For each type profile $\boldsymbol{\tau}=(\tau^{k})_{k\in\mathcal{K}}$, we let $y_{a}(\boldsymbol{\tau})=\sum_{k\in\mathcal{K}}y_{a}^{k}(\tau^{k})$ be the interim total flow on action $a$ and $\boldsymbol{y}(\boldsymbol{\tau})=\left(y_{a}(\boldsymbol{\tau})\right)_{a\in\mathcal{A}}\in\Delta(\mathcal{A})$ be the interim total flow profile, given $\boldsymbol{\tau}$. We let $\operatorname{\mathsf{P}}(\tau^{k})=\sum_{\theta\in\Theta}\sum_{\boldsymbol{\tau}^{-k}\in\mathcal{T}^{-k}}p(\theta)\pi(\tau^{k},\boldsymbol{\tau}^{-k}\mid\theta)$ denote the total probability of type $\tau^{k}$ induced by $p$ and $\pi$, where $\mathcal{T}^{-k}\coloneqq\times_{j\neq k}\mathcal{T}^{j}$. Under the interim flow profile $\widehat{\boldsymbol{y}}$, the payoff of population $k$, given type $\tau^{k}$ and action $a$, such that $y_{a}^{k}(\tau^{k})>0$, is
[…]
The list $(u_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}}))_{k,\tau^{k},a}$ is called an interim payoff profile.
The following definition provides an analog of the usual Bayesian Nash ($\varepsilon$-)equilibrium for nonatomic games.
Definition 2.4.
[…]
$\sum_{\theta\in\Theta}\sum_{\boldsymbol{\tau}^{-k}\in\mathcal{T}^{-k}}\frac{p(\theta)\pi(\tau^{k},\boldsymbol{\tau}^{-k}\mid\theta)}{\operatorname{\mathsf{P}}(\tau^{k})}u_{a}(\boldsymbol{y}(\boldsymbol{\tau}),\theta)\geq\sum_{\theta\in\Theta}\sum_{\boldsymbol{\tau}^{-k}\in\mathcal{T}^{-k}}\frac{p(\theta)\pi(\tau^{k},\boldsymbol{\tau}^{-k}\mid\theta)}{\operatorname{\mathsf{P}}(\tau^{k})}u_{b}(\boldsymbol{y}(\boldsymbol{\tau}),\theta)-\varepsilon.$ (2.5)
That is, for each population $k$ that observes type $\tau^{k}$, a positive mass chooses action $a$ only if, up to $\varepsilon$, it yields the highest conditional expected payoff. Notice that this is equivalent to the following inequality for all $k\in\mathcal{K}$, all $\tau^{k}\in\mathcal{T}^{k}$ and all $a,b\in\mathcal{A}$ such that $y_{a}^{k}(\tau^{k})>0$:
$\sum_{\theta\in\Theta}\sum_{\boldsymbol{\tau}^{-k}\in\mathcal{T}^{-k}}p(\theta)\pi(\tau^{k},\boldsymbol{\tau}^{-k}\mid\theta)u_{a}(\boldsymbol{y}(\boldsymbol{\tau}),\theta)\geq\sum_{\theta\in\Theta}\sum_{\boldsymbol{\tau}^{-k}\in\mathcal{T}^{-k}}p(\theta)\pi(\tau^{k},\boldsymbol{\tau}^{-k}\mid\theta)u_{b}(\boldsymbol{y}(\boldsymbol{\tau}),\theta)-\varepsilon\operatorname{\mathsf{P}}(\tau^{k}).$ (2.6)
A Bayesian Wardrop equilibrium (BWE) is a Bayesian Wardrop $\varepsilon$-equilibrium with $\varepsilon=0$.
An outcome $\mu\colon\Theta\to\Delta(\mathcal{Y})$ is a Bayesian Wardrop $\varepsilon$-equilibrium outcome of $(\Gamma,\mathcal{I})$ if there exists a Bayesian Wardrop $\varepsilon$-equilibrium $\left(\boldsymbol{y}^{k}(\tau^{k})\right)_{k,\tau^{k}}$ such that for every $\theta\in\Theta$ and for every $\boldsymbol{z}\in\mathcal{Y}$,
$\mu(\boldsymbol{z}\mid\theta)=\sum_{\boldsymbol{\tau}\colon\boldsymbol{y}(\boldsymbol{\tau})=\boldsymbol{z}}\pi(\boldsymbol{\tau}\mid\theta).$ (2.7)
We let $\operatorname{\mathsf{BWE_{\varepsilon}}}(\mathcal{I})$ be the set of Bayesian Wardrop $\varepsilon$-equilibrium outcomes of $(\Gamma,\mathcal{I})$.
An information structure is direct if $\mathcal{T}^{k}=\mathcal{A}$ for all $k\in\mathcal{K}$. The associated obedient interim flow profile is the one obtained when population $k$ observes $\tau^{k}=a$, interprets it as an action recommendation, and obeys it by choosing $a$. For every direct information structure $\mathcal{I}$, we let $\breve{\boldsymbol{y}}_{\mathcal{I}}$ denote the obedient interim flow profile, defined as $\breve{\boldsymbol{y}}_{\mathcal{I},a}^{k}(\tau^{k})\coloneqq\gamma^{k}\cdot\operatorname{\mathds{1}}\left\{\tau^{k}=a\right\}$, for all $k\in\mathcal{K},\tau^{k}\in\mathcal{T}^{k},a\in\mathcal{A}$. The associated outcome is denoted by $\breve{\mu}_{\mathcal{I}}$.
3 Examples
The following examples illustrate the notion of BCWE in two simple games and show how to implement the total welfare-maximizing BCWE as an obedient BWE of the game extended with a direct information structure. The examples also show that restricting attention to public information or deterministic flows is not without loss of generality in terms of efficiency.
Example 3.1 (Platform access with network effects and congestion).
In the following game, studied by Mitsche et al., (2013), there is a single state, and the action set is $\mathcal{A}=\left\{a,b\right\}$. The payoff functions are:
$u_{a}(\boldsymbol{y})=1,\quad u_{b}(\boldsymbol{y})=\min\left\{4y_{b},4-4y_{b}\right\}.$ (3.1)
Eq.˜3.1 says that action $b$ generates positive payoff externalities when $y_{b}<1/2$ and negative payoff externalities when $y_{b}>1/2$. This can be interpreted as a platform access decision, where action $b$ corresponds to joining a platform whose value increases with moderate participation due to network effects but deteriorates when participation becomes too high because of congestion, while action $a$ represents a safe outside option with constant payoff 1. This game without information admits three BWE outcomes: $(1,0)$, $(3/4,1/4)$, and $(1/4,3/4)$, all of which yield a total payoff of $1$. Public information induces a mixing over these outcomes, so the expected total payoff remains $1$. Consider now the flows $\boldsymbol{z}=(1,0)$ and $\boldsymbol{w}=(1/2,1/2)$. The outcome $\mu$, defined by
$\mu(\boldsymbol{z})=\frac{1}{3},\quad\mu(\boldsymbol{w})=\frac{2}{3},$ (3.2)
is a BCWE, since it satisfies the equilibrium constraints:
$\begin{split}\frac{2}{3}=\frac{1}{3}z_{a}u_{a}(\boldsymbol{z})+\frac{2}{3}w_{a}u_{a}(\boldsymbol{w})&\geq\frac{1}{3}z_{a}u_{b}(\boldsymbol{z})+\frac{2}{3}w_{a}u_{b}(\boldsymbol{w})=\frac{2}{3},\\ \frac{2}{3}=\frac{1}{3}z_{b}u_{b}(\boldsymbol{z})+\frac{2}{3}w_{b}u_{b}(\boldsymbol{w})&\geq\frac{1}{3}z_{b}u_{a}(\boldsymbol{z})+\frac{2}{3}w_{b}u_{a}(\boldsymbol{w})=\frac{1}{3}.\end{split}$ (3.3)
The total payoff of this BCWE is $4/3$, which is higher than the total payoff with public information. From the platform’s perspective, private recommendations are used to correlate users’ decisions so that the first-best participation level $(1/2,1/2)$ is achieved with probability $2/3$, while the inefficient outcome in which all players choose $b$ never occurs.
The welfare-maximizing BCWE is (partially) implemented as a BWE with the following direct information structure: there are two populations of size $\gamma^{1}=\gamma^{2}=1/2$, and
$\pi(a,a)=\pi(a,b)=\pi(b,a)=\frac{1}{3}.$ (3.4)
Operationally, the platform can achieve this outcome by partitioning users into two equally sized groups and privately recommending access to at most one group at a time.
Example 3.2 (Entry on a market with uncertain production cost).
We consider a game with binary action set $\mathcal{A}=\left\{a,b\right\}$, where action $a$ indicates staying out of the market, and action $b$ indicates entering the market. Firms on the market produce one unit each and sell it at a market price decreasing with total production. The state $\theta$ is the common production cost, which can take values 0 or $1$. This game is a particular instance of the Cournot model described in Section˜5.1.
So we let $\Theta=\left\{0,1\right\}$, and for $q\in(0,1)$,
$p(0)=1-q,\quad\text{and}\quad p(1)=q.$ (3.5)
Let us consider deterministic outcomes, where the flow profile in state $\theta$ is some $\boldsymbol{y}(\theta)=\left(y_{a}(\theta),y_{b}(\theta)\right)=\left(1-y_{b}(\theta),y_{b}(\theta)\right)\in\mathcal{Y}$. The inverse demand depends linearly on total production and is given by
$\rho(y_{b}(\theta))=1-y_{b}(\theta).$ (3.6)
Firms maximize profits, which are
$\begin{split}u_{a}(\boldsymbol{y},\theta)&=0,\\ u_{b}(\boldsymbol{y},\theta)&=\rho(y_{b}(\theta))-\theta.\end{split}$ (3.7)
The total profit in state $\theta$ is
$y_{a}(\theta)u_{a}(\boldsymbol{y}(\theta),\theta)+y_{b}(\theta)u_{b}(\boldsymbol{y}(\theta),\theta)=y_{b}(\theta)(1-y_{b}(\theta)-\theta).$ (3.8)
The expected total profit is
$(1-q)\,y_{b}(0)(1-y_{b}(0))-q\,y_{b}(1)y_{b}(1).$ (3.9)
The maximizer over $0,1]^{2}$ is $\left(y_{b}(0)=1/2,y_{b}(1)=0\right)$, with expected total profit equal to (1q)/4(1-q)/4.
For deterministic outcomes, the BCWE equilibrium constraints reduce to:
$(1-q)(1-y_{b}(0))^{2}$ $\leq q\,y_{b}(1)(1-y_{b}(1)),$ (3.10)
$(1-q)y_{b}(0)(1-y_{b}(0))$ $\geq q\,(y_{b}(1))^{2}.$ (3.11)
Maximizing expected total profit subject to Eqs.˜3.11 and 3.10 yields a unique solution with (3.10) binding and (3.11) slack:
$y_{b}^{*}(0)=1-\frac{\sqrt{q}}{2},\quad y_{b}^{*}(1)=\frac{1-\sqrt{q}}{2}.$ (3.12)
The associated expected total profit is $\left(\sqrt{q}-q\right)/2$, which is strictly smaller than the unconstrained benchmark but strictly larger than the expected total profit achievable under public information, i.e., 0. By Proposition˜6.1, restricting attention to deterministic flows is without loss of optimality in this binary-action environment, so $(y_{b}^{*}(0),y_{b}^{*}(1))$ is the total profit-maximizing BCWE outcome.
Similarly to the previous example, as long as $\sqrt{q}$ is a rational number, this deterministic BCWE can be implemented as an obedient BWE using a direct information structure with finitely many populations, by recommending action $b$ to a fraction $y_{b}^{*}(\theta)$ of firms in each state $\theta$. For example, if $q=1/4$, then
$y_{b}^{*}(0)=1-\frac{\sqrt{q}}{2}=\frac{3}{4},\qquad y_{b}^{*}(1)=\frac{1-\sqrt{q}}{2}=\frac{1}{4}.$ (3.13)
Consider four populations of equal size $\gamma^{k}=1/4$ for every $k$, and a direct information structure such that, in state 0, exactly three populations are recommended action $b$ and one population is recommended action $a$, whereas in state $1$ exactly one population is recommended action $b$ and three populations are recommended action $a$ (uniformly over the choice of the recommended populations). This induces the deterministic outcome $(y_{b}(0),y_{b}(1))=(3/4,1/4)$. As shown in Section˜5, this total profit-maximizing outcome is fully implemented in this game because it is a concave potential game. Moreover, when $\sqrt{q}$ is not rational, the total profit-maximizing outcome can be implemented with arbitrary precision by using a sufficiently large number of populations.
4 Partial Implementation
It is immediate that any BWE outcome, for any possible information structure (direct or indirect), is in the set of BCWE. This simple observation explicitly illustrates how to construct some BCWE: choose an information structure $\left(\boldsymbol{\gamma},\mathcal{T},\pi\right)$ for disclosing information, and choose a BWE of the basic game extended with this information structure.
The next theorem establishes the other direction of the revelation principle (Myerson,, 1982, Bergemann and Morris,, 2016), at least approximately. Specifically, it shows that if $\mu$ is a BCWE with finite support and rational flows, then there exists a direct information structure such that $\mu$ is implemented as an obedient BWE outcome. If the support of $\mu$ is not finite or if some flows in the support are not rational, then $\mu$ can be approximately implemented as a Bayesian Wardrop $\varepsilon$-equilibrium outcome for every $\varepsilon>0$.
Theorem 4.1.
Let $\mu$ be a BCWE and $\psi\colon\mathcal{Y}\times\Theta\to\mathbb{R}$ be a continuous function. For all $\varepsilon>0$, there exists a direct information structure $\mathcal{I}$ such that the obedient interim flow profile $\breve{\boldsymbol{y}}_{\mathcal{I}}$ is a Bayesian Wardrop $\varepsilon$-equilibrium of $(\Gamma,\mathcal{I})$ and
$\left\lvert\int\psi(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)-\int\psi(\boldsymbol{y},\theta)\ \textup{d}\breve{\mu}_{\mathcal{I}}(\boldsymbol{y}\mid\theta)\right\rvert\leq\varepsilon,\quad\forall\theta\in\Theta.$ (4.1)
Further, if $\mu$ has finite support and if all flows in the support of $\mu$ are rational, then there exists an information structure $\mathcal{I}$ such that $\varepsilon=0$ and $\mu=\breve{\mu}_{\mathcal{I}}$.
In Theorem˜4.1, the function $\psi$ represents any possible designer welfare function, and Eq.˜4.1 means that for every state, the interim expected payoff of the designer under outcome $\breve{\mu}_{\mathcal{I}}$ is $\varepsilon$-close to the interim expected payoff under the BCWE $\mu$. The proof of the theorem also shows that, for every $L>0$, the same $\mathcal{I}$ can be chosen for all $L$-Lipschitz functions $\psi$.
Proof of Theorem˜4.1.
Koessler et al., (2025, Lemma 2) prove that BCWE with finite support are dense within the set of Bayes correlated Wardrop equilibria. So, it is enough to prove the result for $\mu$ with finite support. Fix $\varepsilon>0$ and a BCWE $\mu$ with finite support $\mathcal{Y}^{\ast}\subset\mathcal{Y}$. Let $\mathcal{A}^{*}$ be the set of actions $a$ actually used under $\mu$, i.e., such that
$\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}p(\theta)\mu(\boldsymbol{y}\mid\theta)y_{a}>0.$ (4.2)
Let then
$\varepsilon_{0}=\min\Big\{\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}p(\theta)\mu(\boldsymbol{y}\mid\theta)y_{a}\text{ s.t. }a\in\mathcal{A}^{*}\Big\}>0.$ (4.3)
Fix an arbitrary $\eta>0$ and approximate flows in $\mathcal{Y}^{\ast}$ by vectors with rational components as follows. There exists $K\in\mathbb{N}_{+}$, such that for every $\boldsymbol{y}\in\mathcal{Y}^{\ast}$ and every $a\in\mathcal{A}^{*}$, there exists $N(y_{a})\in\mathbb{N}$, such that,
$\sum_{a\in\mathcal{A}^{*}}N(y_{a})=K\quad\text{and}\quad\forall a\in\mathcal{A}^{*},\left\lvert\frac{N(y_{a})}{K}-y_{a}\right\rvert\leq\eta.$ (4.4)
For each $\boldsymbol{y}\in\mathcal{Y}^{\ast}$, let $\widetilde{\boldsymbol{y}}$ denote the flow profile $(\widetilde{y}_{a})_{a\in\mathcal{A}^{*}}$ with $\widetilde{y}_{a}=N(y_{a})/K$.
We construct a direct information structure $\mathcal{I}$ as follows. There are $|\mathcal{K}|=K$ populations and each population $k$ has the same size $\gamma^{k}=1/K$. For each population $k$, the set of types is $\mathcal{T}^{k}=\mathcal{A}^{*}$. Conditionally on state $\theta$, a flow profile $\boldsymbol{y}\in\mathcal{Y}^{\ast}$ is drawn with probability $\mu(\boldsymbol{y}\mid\theta)$, then action $a$ is recommended to a subset of $N(y_{a})$ populations chosen uniformly, so that, conditionally on $\boldsymbol{y}$ and $\theta$, the probability that population $k$ is recommended $a$ is $\widetilde{y}_{a}$. We let $\breve{\mu}_{\mathcal{I}}$ be the induced obedient outcome which puts probability $\mu(\boldsymbol{y}\mid\theta)$ on $\widetilde{\boldsymbol{y}}$ in state $\theta$.
For each population $k$, the conditional expected payoff of playing $b$ when $a$ is recommended is
$\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}\frac{p(\theta)\mu(\boldsymbol{y}\mid\theta)\widetilde{y}_{a}}{\operatorname{\mathsf{Q}}(a)}u_{b}(\widetilde{\boldsymbol{y}},\theta),$ (4.5)
with
$\operatorname{\mathsf{Q}}(a)=\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}p(\theta)\mu(\boldsymbol{y}\mid\theta)\widetilde{y}_{a}.$ (4.6)
For each pair of actions $(a,b)$, the mapping $\boldsymbol{y}\mapsto y_{a}u_{b}(\boldsymbol{y},\theta)$ is uniformly continuous on the compact set $\mathcal{Y}$. Thus, there exists a common modulus of continuity $\omega(\,\cdot\,)$ with $\lim_{\eta\searrow 0}\omega(\eta)=0$ such that, for all $\theta\in\Theta$, for all $(a,b)\in\mathcal{A}\times\mathcal{A}$, for all $(\boldsymbol{y},\boldsymbol{z})\in\mathcal{Y}\times\mathcal{Y}$,
[…]
Since $\mu$ is BCWE, for all $a,b\in\mathcal{A}$,
$\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}p(\theta)\mu(\boldsymbol{y}\mid\theta)y_{a}u_{a}(\boldsymbol{y},\theta)\geq\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}p(\theta)\mu(\boldsymbol{y}\mid\theta)y_{a}u_{b}(\boldsymbol{y},\theta).$ (4.8)
From Eqs.˜4.8, 4.4 and 4.7 it follows that
$\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}p(\theta)\mu(\boldsymbol{y}\mid\theta)\widetilde{y}_{a}u_{a}(\widetilde{\boldsymbol{y}},\theta)\geq\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}p(\theta)\mu(\boldsymbol{y}\mid\theta)\widetilde{y}_{a}u_{b}(\widetilde{\boldsymbol{y}},\theta)-2\omega(\eta).$ (4.9)
Thus
$\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}\frac{p(\theta)\mu(\boldsymbol{y}\mid\theta)\widetilde{y}_{a}}{\operatorname{\mathsf{Q}}(a)}u_{a}(\widetilde{\boldsymbol{y}},\theta)\geq\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}\frac{p(\theta)\mu(\boldsymbol{y}\mid\theta)\widetilde{y}_{a}}{\operatorname{\mathsf{Q}}(a)}u_{b}(\widetilde{\boldsymbol{y}},\theta)-\frac{2\omega(\eta)}{\operatorname{\mathsf{Q}}(a)}.$ (4.10)
From Eq.˜4.4 we have
$\operatorname{\mathsf{Q}}(a)=\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}p(\theta)\mu(\boldsymbol{y}\mid\theta)\widetilde{y}_{a}\geq\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}p(\theta)\mu(\boldsymbol{y}\mid\theta)(y_{a}-\eta)\geq\varepsilon_{0}-\eta\geq\frac{\varepsilon_{0}}{2},$ (4.11)
for $\eta\leq\varepsilon_{0}/2$. It follows that
$\frac{2\omega(\eta)}{\operatorname{\mathsf{Q}}(a)}\leq\frac{2\omega(\eta)}{\varepsilon_{0}/2}.$ (4.12)
[…]
Now, take a continuous $\psi\colon\mathcal{Y}\times\Theta\to\mathbb{R}$. Since $\mathcal{Y}\times\Theta$ is compact, $\psi$ is uniformly continuous and admits a modulus of continuity $\omega_{\psi}$ such that for all $\theta\in\Theta$, for all $(\boldsymbol{y},\boldsymbol{z})\in\mathcal{Y}\times\mathcal{Y}$,
[…]
For each $\theta\in\Theta$, we have
$\begin{split}\left\lvert\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}\mu(\boldsymbol{y}\mid\theta)\psi(\widetilde{\boldsymbol{y}},\theta)-\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}\mu(\boldsymbol{y}\mid\theta)\psi(\boldsymbol{y},\theta)\right\rvert&\leq\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}\mu(\boldsymbol{y}\mid\theta)\left\lvert\psi(\widetilde{\boldsymbol{y}},\theta)-\psi(\boldsymbol{y},\theta)\right\rvert\\ &\leq\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}\mu(\boldsymbol{y}\mid\theta)\omega_{\psi}(\eta)\\ &\leq\omega_{\psi}(\eta),\end{split}$ (4.14)
which is less or equal to $\varepsilon$ for $\eta$ small.
If flows in $\mathcal{Y}^{\ast}$ have rational coefficients with common denominator $K$, then we can choose $\widetilde{\boldsymbol{y}}=\boldsymbol{y}$ and $\eta=0$, and we have a Bayesian Wardrop 0-equilibrium. Finally, notice that if $\psi$ is $L$-Lipschitz (for the sup norm on flows), then we can choose $\omega_{\psi}(\eta)=L\eta$. Thus, for given $\varepsilon$, the choice of $\eta$ and $\mathcal{I}$ is the same across all such $\psi$. ∎
5 Potential Games and Full Implementation
In this section, we study full implementation of BCWE in games with a concave potential.
5.1 Potential Games
Definition 5.1.
A basic game $\Gamma=\left(\mathcal{A},\Theta,p,\boldsymbol{u}\right)$ is called a potential game if for every $\theta\in\Theta$, there exists an open superset $\widetilde{\mathcal{Y}}$ of $\mathcal{Y}$ in $\mathbb{R}^{\mathcal{A}}$ and a continuously differentiable function $\Phi_{\theta}\colon\widetilde{\mathcal{Y}}\to\mathbb{R}$ such that, for every $a\in\mathcal{A}$ and $\boldsymbol{y}\in\mathcal{Y}$,
$\frac{\partial\Phi_{\theta}(\boldsymbol{y})}{\partial y_{a}}=u_{a}(\boldsymbol{y},\theta).$ (5.1)
The function $\Phi_{\theta}$ is called the potential in state $\theta$.
As a first example, consider symmetric congestion games, which are defined by: a finite set of resources $\mathcal{E}$; a set of actions $\mathcal{A}\subseteq 2^{\mathcal{E}}$; for each resource $e\in\mathcal{E}$, a continuous cost function $c_{e}\colon\mathbb{R}_{+}\times\Theta\to\mathbb{R}_{+}$. The payoff of choosing action $a\in\mathcal{A}$ in state $\theta$ is the negative of the sum of the costs of all resources used in $a$,
$u_{a}(\boldsymbol{y},\theta)\coloneqq-\sum_{e\in a}c_{e}(x_{e},\theta),$ (5.2)
where $x_{e}\coloneqq\sum_{a\in\mathcal{A},a\ni e}y_{a}$ is the load on resource $e$ induced by the flow $\boldsymbol{y}$. The potential function of a congestion game is
$\Phi_{\theta}(\boldsymbol{y})=-\sum_{e\in\mathcal{E}}\int_{0}^{x_{e}}c_{e}(t,\theta)\ \textup{d}t.$ (5.3)
When the cost function of each resource is increasing for every state, the potential is concave. In singleton congestion games each action uses a single resource, i.e., $\mathcal{A}=\{\{e\},e\in\mathcal{E}\}$. In this class of games, if the cost functions are strictly increasing, the potential is strictly concave.
The following lemma shows that, unlike what happens in games with a finite set of players (see Monderer and Shapley,, 1996), not all potential games are congestion games.
Lemma 5.2.
Fix an arbitrary state $\theta$. A game with a potential $\Phi_{\theta}\in C^{3}$ is a congestion game only if
$\frac{\partial^{3}\Phi_{\theta}(\boldsymbol{y})}{\partial y_{a}^{2}\partial y_{b}}=\frac{\partial^{3}\Phi_{\theta}(\boldsymbol{y})}{\partial y_{b}^{2}\partial y_{a}}.$ (5.4)
Proof.
Consider a potential game with a potential function $\Phi_{\theta}\in C^{3}$, and suppose that it is a congestion game, with cost functions $c_{e}(\,\cdot\,,\theta)$, which are then twice continuously differentiable. Combine Eqs.˜5.1 and 5.2 to get,
$\frac{\partial\Phi_{\theta}(\boldsymbol{y})}{\partial y_{a}}=-\sum_{e\in a}c_{e}(x_{e},\theta).$ (5.5)
This implies that
$\frac{\partial^{2}\Phi_{\theta}(\boldsymbol{y})}{\partial y_{a}\partial y_{b}}=-\sum_{e\in a\cap b}c^{\prime}_{e}(x_{e},\theta)=\frac{\partial^{2}\Phi_{\theta}(\boldsymbol{y})}{\partial y_{b}\partial y_{a}}.$ (5.6)
Observe that $x_{e}=\sum_{a^{\prime}\in\mathcal{A},a\ni e}y_{a^{\prime}}$. Because we consider only resources $e\in a\cap b$, the flow $y_{a}$ or $y_{b}$ only appears in the sum $y_{a}+y_{b}$. Therefore, taking derivative once more with respect to either $y_{a}$ or $y_{b}$ yields then the same result. ∎
To go beyond congestion games, we consider the class of aggregative games. This class was implicitly introduced by Dubey et al., (1980), formalized by Corchón, (1994), and recently studied by Lahkar, (2017) for nonatomic players and a finite set of actions. Aggregative games are also given as examples of potential games in Sandholm, (2010, chapter 3).
An aggregative game is defined as follows. Let the action set $\mathcal{A}$ be a finite subset of $\mathbb{R}_{+}$, so actions can be seen as quantities. For $y\in\Delta(\mathcal{A})$, let the aggregate quantity be $Q(\boldsymbol{y})=\sum_{a\in\mathcal{A}}ay_{a}$. Each payoff $u_{a}$ depends only on the chosen action $a$ and on the aggregate quantity, i.e., for each $a\in\mathcal{A}$ and for each $\theta\in\Theta$, there exists a function $g_{a}(\,\cdot\,,\theta)$ such that $u_{a}(y,\theta)=g_{a}(Q(\boldsymbol{y}),\theta)$.
Consider payoffs of the following form:
$\forall a\in\mathcal{A},u_{a}(y,\theta)=a\rho(Q(\boldsymbol{y}),\theta)-C_{a}(\theta),$ (5.7)
where $\rho(\,\cdot\,,\theta)$ is a function of the aggregate quantity $Q(\boldsymbol{y})$, and $C_{a}(\theta)$ is a number that does not depend on $\boldsymbol{y}$. This game admits the potential:
$\Phi_{\theta}(\boldsymbol{y})=\int_{0}^{Q(\boldsymbol{y})}\rho(t,\theta)\ \textup{d}t-\sum_{a}C_{a}(\theta)y_{a}.$ (5.8)
If $\rho(\,\cdot\,,\theta)$ is a decreasing function of the quantity, then the potential is concave. [^6] Sandholm, (2010) and Lahkar, (2017) also consider the case of increasing $\rho(\,\cdot\,,\theta)$, which models public good contribution games where the value of the good is increasing with the total contribution, and players incur contribution costs. In this case, the potential is convex. This potential can be interpreted as a demand function and the game is a Cournot game, where $C_{a}(\theta)$ is the production cost of quantity $a$. In this model, both the demand and the production costs may be uncertain. In Example˜3.2 only production costs are uncertain and the potential is strictly concave.
With the help of Lemma˜5.2, assuming that $\rho(\,\cdot\,,\theta)$ is twice continuously differentiable, we can check whether the game given by Eq.˜5.7 is a congestion game or not. Suppose that $\rho^{\prime\prime}(\,\cdot\,,\theta)>0$ and assume that there exist two distinct actions $a,b\in\mathcal{A}\setminus\{0\}$. Then
$\frac{\partial^{3}\Phi_{\theta}(\boldsymbol{y})}{\partial y_{a}^{2}\partial y_{b}}=a^{2}b\rho^{\prime\prime}(Q(\boldsymbol{y}),\theta)\neq ab^{2}\rho^{\prime\prime}(Q(\boldsymbol{y}),\theta)=\frac{\partial^{3}\Phi_{\theta}(\boldsymbol{y})}{\partial y_{b}^{2}\partial y_{a}}.$ (5.9)
Therefore, the game is not a congestion game. Take for instance a demand function given by $\rho(q,\theta)=\kappa_{\theta}/(\lambda_{\theta}+q)$.
Proposition 5.3.
A potential game with a potential $\Phi_{\theta}\in C^{2}$ is aggregative if and only if Eq.˜5.7 holds.
Proof.
Take an aggregative game with payoffs $u_{a}(y,\theta)=g_{a}(Q(\boldsymbol{y}),\theta)$ and suppose it admits a potential $\Phi_{\theta}$. Then,
$\frac{\partial\Phi_{\theta}(\boldsymbol{y})}{\partial y_{a}}$ $=g_{a}(Q(\boldsymbol{y}),\theta)$ (5.10)
and, for all $a,b\in\mathcal{A}$,
$\frac{\partial^{2}\Phi_{\theta}(\boldsymbol{y})}{\partial y_{b}\partial y_{a}}$ $=bg^{\prime}_{a}(Q(\boldsymbol{y}),\theta)=ag^{\prime}_{b}(Q(\boldsymbol{y}),\theta)=\frac{\partial^{2}\Phi_{\theta}(\boldsymbol{y})}{\partial y_{a}\partial y_{b}}.$ (5.11)
So for any two actions $a,b\neq 0$, the two ratios
$\frac{g^{\prime}_{a}(Q(\boldsymbol{y}),\theta)}{a}=\frac{g^{\prime}_{b}(Q(\boldsymbol{y}),\theta)}{b}$ (5.12)
are equal to some common function of $Q(\boldsymbol{y})$. Call this function $\rho^{\prime}(Q(\boldsymbol{y}),\theta)$. Eq.˜5.7 then follows by integration, where $C_{a}(\theta)$ is an integration constant. Note that for action 0, Eq.˜5.11 implies that $g^{\prime}_{0}(Q(\boldsymbol{y}),\theta)=0$, so Eq.˜5.7 is valid for $a=0$ as well. ∎
Finally, notice that every game with two actions is a potential game, though the potential need not be concave (see, e.g., Example˜3.1).
5.2 Full Implementation
The next proposition shows that for every potential game with a (strictly) concave potential, the BWE payoff profile (outcome) of the game extended by an information structure is unique.
Proposition 5.4.
Consider a basic game $\Gamma$. If the potential is concave for each state, then for every information structure $\mathcal{I}$, all BWE of $(\Gamma,\mathcal{I})$ have the same payoff profile. If the potential is strictly concave for each state, then for every information structure $\mathcal{I}$ there exists a unique BWE outcome of $(\Gamma,\mathcal{I})$.
To prove Proposition˜5.4, we first show that a BWE of $(\Gamma,\mathcal{I})$ is equivalent to a Wardrop equilibrium of an auxiliary game with complete information and multiple populations, which is also a potential game. If the basic game has a concave potential in each state, then the potential of the auxiliary game is also concave. Uniqueness of Wardrop equilibrium payoffs in the auxiliary game follows from Koessler et al., (2025, Proposition 4). As a consequence, we get uniqueness of BWE payoff profiles for $(\Gamma,\mathcal{I})$. If the basic game has a strictly concave potential in each state, then the potential of the auxiliary game is a strictly concave function of the interim total flow profiles $(\boldsymbol{y}(\boldsymbol{\tau}))_{\boldsymbol{\tau}}$. Uniqueness of the equilibrium interim total flow profile follows from the fact that the Wardrop equilibrium of the auxiliary multi-population game with complete information are the maximizers of the potential (see Sandholm,, 2001, proposition 3.1).
Proof of Proposition˜5.4.
Consider a basic $\Gamma$ and an information structure $\mathcal{I}=\left(\boldsymbol{\gamma},\mathcal{T},\pi\right)$. Define a multi-population auxiliary game with complete information where populations are indexed by $(k,\tau^{k})$, for all $k\in\mathcal{K},\tau^{k}\in\mathcal{T}^{k}$. A flow profile for population $(k,\tau^{k})$ is $\boldsymbol{y}^{k}(\tau^{k})=\left(y_{a}^{k}(\tau^{k})\right)_{a\in\mathcal{A}}\in\Delta_{\gamma^{k}}(\mathcal{A})$. Given the interim flow profile $\widehat{\boldsymbol{y}}=\left(y^{k}_{a}(\tau^{k})\right)_{k,\tau^{k},a}$, the payoff function of population $(k,\tau^{k})$ is
[…]
A Wardrop equilibrium of this auxiliary game is identical to a Bayesian Wardrop equilibrium of $(\Gamma,\mathcal{I})$.
Suppose that in each state $\theta$ there is a potential $\Phi_{\theta}$. Then, let
$\Phi_{\pi}(\widehat{\boldsymbol{y}})\coloneqq\sum_{\theta,\tau}p(\theta)\pi(\boldsymbol{\tau}\mid\theta)\Phi_{\theta}\left(\left(\sum_{k}y_{a}^{k}(\tau^{k})\right)_{a\in\mathcal{A}}\right)=\sum_{\theta,\tau}p(\theta)\pi(\boldsymbol{\tau}\mid\theta)\Phi_{\theta}(\boldsymbol{y}(\boldsymbol{\tau})),$ (5.14)
where we use the notation $y_{a}(\boldsymbol{\tau})=\sum_{k\in\mathcal{K}}y_{a}^{k}(\tau^{k})$ and $\boldsymbol{y}(\boldsymbol{\tau})=\left(y_{a}(\boldsymbol{\tau})\right)_{a\in\mathcal{A}}.$ The function $\Phi_{\pi}$ is a potential for the auxiliary game and
[…]
If $\Phi_{\theta}$ is strictly concave for each $\theta$, then $\Phi_{\pi}$ is a strictly concave function of the vector $(\boldsymbol{y}(\boldsymbol{\tau}))_{\boldsymbol{\tau}}$. By Sandholm, (2001, Proposition 3.1), the equilibrium flows are the maximizers of the potential function. Thus, there is a unique equilibrium vector $(\boldsymbol{y}(\boldsymbol{\tau}))_{\boldsymbol{\tau}}$, and therefore a unique equilibrium outcome, given by (2.7). If $\Phi_{\pi}$ is (weakly) concave, uniqueness of the equilibrium payoff profile $(U_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}}))_{k,\tau^{k},a}$ follows from Koessler et al., (2025, proposition 4). Uniqueness of the BWE payoff profile follows from Eq.˜5.13. ∎
The next theorem is a full implementation result for games with a strictly concave potential. It shows that if $\mu$ is a BCWE with finite support and rational flows, then there exists a direct information structure such that $\mu$ is implemented as a unique Bayesian Wardrop equilibrium outcome. Furthermore, the information structure is direct and the BWE is obedient. If the support of $\mu$ is not finite or if some flows in the support are not rational, then $\mu$ is close to the unique BWE outcome $\mu^{\ast}$ for some direct information structure $\mathcal{I}$.
Theorem 5.5.
Let $\Gamma$ be a basic game with a strictly concave potential in each state. Let $\mu$ be a BCWE and let $\psi\colon\mathcal{Y}\times\Theta\to\mathbb{R}$ be a continuous function. For all $\varepsilon>0$, there exists a direct information structure $\mathcal{I}$ such that the Bayesian Wardrop equilibrium outcome $\mu^{\ast}$ of $(\Gamma,\mathcal{I})$ is unique and
$\left\lvert\int\psi(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)-\int\psi(\boldsymbol{y},\theta)\ \textup{d}\mu^{\ast}(\boldsymbol{y}\mid\theta)\right\rvert\leq\varepsilon,\quad\forall\theta\in\Theta.$ (5.16)
In addition, if $\mu$ has finite support with rational flows, then we can choose $\varepsilon=0$ and the unique Bayesian Wardrop equilibrium outcome $\mu^{\ast}$ is the obedient outcome $\breve{\mu}_{\mathcal{I}}$.
As before, $\psi$ represents any possible welfare function for the designer. Also, for every $L>0$, we can choose the same $\mathcal{I}$ for all $L$-Lipschitz $\psi$.
The structure of the proof is as follows. Consider the construction given in the proof of Theorem˜4.1: Approximate $\mu$ by a distribution over a finite set of rational flows and construct a direct information structure such that the obedient outcome is a Bayesian Wardrop $\alpha$-equilibrium for some $\alpha>0$, and is close to $\mu$. The difficulty is that the obedient outcome need not be an exact Bayesian Wardrop equilibrium and that we don’t know how to construct this BWE explicitly. We show that any Bayesian Wardrop $\alpha$-equilibrium is an $\alpha$-maximizer of the potential of the game with the constructed information structure, and that all $\alpha$-maximizers are $\varepsilon$-close to each other for some $\varepsilon>0$. In particular, the unique BWE outcome $\mu^{\ast}$ is close to the obedient outcome and thus, also close to $\mu$. This reasoning is summarized in the following proposition.
Proposition 5.6.
Let $\Gamma$ be a basic game with a strictly concave potential in each state. Let $\mu$ be a BCWE and let $\psi\colon\mathcal{Y}\times\Theta\to\mathbb{R}$ be a continuous function. For all $\varepsilon>0$, there exist $\alpha>0$ and a direct information structure $\mathcal{I}$ such that the Bayesian Wardrop equilibrium outcome $\mu^{\ast}$ of $(\Gamma,\mathcal{I})$ is unique, and for all $\widetilde{\mu}\in\operatorname{\mathsf{BWE_{\alpha}}}(\mathcal{I})$,
$\left\lvert\int\psi(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)-\int\psi(\boldsymbol{y},\theta)\ \textup{d}\widetilde{\mu}(\boldsymbol{y}\mid\theta)\right\rvert\leq\varepsilon,\quad\forall\theta\in\Theta.$ (5.17)
Note that $\mu^{\ast}$ belongs to $\operatorname{\mathsf{BWE_{\alpha}}}(\mathcal{I})$ for all $\alpha\geq 0$.
Proof of Proposition˜5.6.
We take up the construction from the proof of Theorem˜4.1: fix $\alpha>0$, assume that $\mu$ has finite support, and consider the information structure $\mathcal{I}$ constructed there with parameter $\eta>0$, such that the obedient flow profile $\breve{\boldsymbol{y}}_{\mathcal{I}}$ with outcome $\breve{\mu}_{\mathcal{I}}$ is a Bayesian Wardrop $\alpha$-equilibrium and
$\left\lvert\int\psi(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)-\int\psi(\boldsymbol{y},\theta)\ \textup{d}\breve{\mu}_{\mathcal{I}}(\boldsymbol{y}\mid\theta)\right\rvert\leq\alpha,\quad\forall\theta\in\Theta.$ (5.18)
Claim 5.7.
Suppose that the potential is (weakly) concave in each state. For all $\alpha>0$, any Bayesian Wardrop $\alpha$-equilibrium of the game with information structure $\mathcal{I}$, is an $\alpha$-maximizer of the associated potential.
Proof of ˜5.7.
Consider an arbitrary Bayesian Wardrop $\alpha$-equilibrium $\widehat{\boldsymbol{y}}$. For all $k\in\mathcal{K}$, for all $\tau^{k}\in\mathcal{T}^{k}$, and for all $a,b\in\mathcal{A}$, if $y_{a}^{k}(\tau^{k})>0$ then,
$U_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}})\geq U_{b}^{k,\tau^{k}}(\widehat{\boldsymbol{y}})-\alpha\operatorname{\mathsf{P}}(\tau^{k})\geq U_{b}^{k,\tau^{k}}(\widehat{\boldsymbol{y}})-\alpha,$ (5.19)
where $U_{a}^{k,\tau^{k}}$ is defined as in (5.13). This implies that, for all $k\in\mathcal{K}$, for all $\tau^{k}\in\mathcal{T}^{k}$,
$\sum_{a\in\mathcal{A}}\frac{y_{a}^{k}(\tau^{k})}{\gamma^{k}}U_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}})\geq\max_{b\in\mathcal{A}}U_{b}^{k,\tau^{k}}(\widehat{\boldsymbol{y}})-\alpha=\max_{\widehat{\boldsymbol{z}}}\sum_{a\in\mathcal{A}}\frac{z_{a}^{k}(\tau^{k})}{\gamma^{k}}U_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}})-\alpha.$ (5.20)
That is, for all $k\in\mathcal{K}$, for all $\tau^{k}\in\mathcal{T}^{k}$ and all $\widehat{\boldsymbol{z}}$,
$\sum_{a\in\mathcal{A}}\left(y_{a}^{k}(\tau^{k})-z_{a}^{k}(\tau^{k})\right)U_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}})\geq-\gamma^{k}\alpha.$ (5.21)
Summing up over $k$ and $\tau^{k}$, we get,
$\left\langle\nabla\Phi_{\pi}(\widehat{\boldsymbol{y}}),\widehat{\boldsymbol{y}}-\widehat{\boldsymbol{z}}\right\rangle=\sum_{k}\sum_{\tau^{k}}\sum_{a\in\mathcal{A}}\left(y_{a}^{k}(\tau^{k})-z_{a}^{k}(\tau^{k})\right)U_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}})\geq-\sum_{k}\gamma^{k}\alpha=-\alpha,$ (5.22)
and since $\Phi_{\pi}$ is concave, this gives
$\Phi_{\pi}(\widehat{\boldsymbol{y}})-\Phi_{\pi}(\widehat{\boldsymbol{z}})\geq\left\langle\nabla\Phi_{\pi}(\widehat{\boldsymbol{y}}),\widehat{\boldsymbol{y}}-\widehat{\boldsymbol{z}}\right\rangle\geq-\alpha.$ (5.23)
Thus $\Phi_{\pi}(\widehat{\boldsymbol{y}})\geq\max\Phi_{\pi}-\alpha$. That is, any Bayesian Wardrop $\alpha$-equilibrium is an $\alpha$-maximizer of the potential. ∎
From strict concavity, let $(\boldsymbol{y}^{*}(\boldsymbol{\tau}))_{\boldsymbol{\tau}}$ be the unique Bayesian Wardrop equilibrium flow vector. We argue now that for any $\alpha$-maximizer $\widehat{\boldsymbol{y}}$ of $\Phi_{\pi}$, the flow vector $(\boldsymbol{y}(\boldsymbol{\tau}))_{\boldsymbol{\tau}}$ is close to $(\boldsymbol{y}^{*}(\boldsymbol{\tau}))_{\boldsymbol{\tau}}$.
Claim 5.8.
For all $\delta>0$, there exists $\alpha>0$ such that for all $\widehat{\boldsymbol{y}}$, $\Phi_{\pi}(\widehat{\boldsymbol{y}})\geq\max\Phi_{\pi}-\alpha$ implies that for all $a$ and all $\boldsymbol{\tau}$, $\left\lvert y_{a}(\boldsymbol{\tau})-y^{*}_{a}(\boldsymbol{\tau})\right\rvert\leq\delta$.
Proof of ˜5.8.
By contradiction, suppose that there exists $\delta>0$ such that for all $\alpha=1/n$ with $n\in\mathbb{N}$, there exists $\widehat{\boldsymbol{y}}_{n}$ with $\Phi_{\pi}(\widehat{\boldsymbol{y}}_{n})\geq\max\Phi_{\pi}-1/n$ and $\max_{a,\boldsymbol{\tau}}\left\lvert y_{n,a}(\boldsymbol{\tau})-y^{*}_{a}(\boldsymbol{\tau})\right\rvert>\delta$. By compactness, take a subsequence and assume that $\widehat{\boldsymbol{y}}_{n}\to\widehat{\boldsymbol{y}}_{\infty}$ as $n\to\infty$. Letting $n\to\infty$, we get that $\Phi_{\pi}(\widehat{\boldsymbol{y}}_{\infty})\geq\max\Phi_{\pi}$ and $\max_{a,\boldsymbol{\tau}}\left\lvert\widehat{y}_{\infty,a}(\boldsymbol{\tau})-y^{*}_{a}(\boldsymbol{\tau})\right\rvert\geq\delta$. This contradicts the uniqueness of $y^{*}_{a}(\boldsymbol{\tau})$. ∎
Take a Bayesian Wardrop $\alpha$-equilibrium with flow vector $(\widetilde{\boldsymbol{y}}(\boldsymbol{\tau}))_{\boldsymbol{\tau}}$ and outcome $\widetilde{\mu}$. We have for each $\theta\in\Theta$,
$\begin{split}\left\lvert\int\psi(\boldsymbol{y},\theta)\ \textup{d}\mu^{\ast}(\boldsymbol{y}\mid\theta)-\int\psi(\boldsymbol{y},\theta)\ \textup{d}\widetilde{\mu}(\boldsymbol{y}\mid\theta)\right\rvert&=\left\lvert\sum_{\boldsymbol{\tau}}\pi(\boldsymbol{\tau}\mid\theta)\psi(\boldsymbol{y}^{*}(\boldsymbol{\tau}),\theta)-\sum_{\boldsymbol{\tau}}\pi(\boldsymbol{\tau}\mid\theta)\psi(\widetilde{\boldsymbol{y}}(\boldsymbol{\tau}),\theta)\right\rvert\\ &\leq\sum_{\boldsymbol{\tau}}\pi(\boldsymbol{\tau}\mid\theta)\left\lvert\psi(\boldsymbol{y}^{*}(\boldsymbol{\tau}),\theta)-\psi(\widetilde{\boldsymbol{y}}(\boldsymbol{\tau}),\theta)\right\rvert\\ &\leq\omega_{\psi}(\delta).\end{split}$
Summing up, for all $\theta\in\Theta$, we have the following inequalities. By construction, we have
$\left\lvert\int\psi(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)-\int\psi(\boldsymbol{y},\theta)\ \textup{d}\breve{\mu}_{\mathcal{I}}(\boldsymbol{y}\mid\theta)\right\rvert\leq\alpha.$ (5.24)
Moreover, because the obedient interim profile is a Bayesian Wardrop $\alpha$-equilibrium, we have
$\left\lvert\int\psi(\boldsymbol{y},\theta)\ \textup{d}\breve{\mu}_{\mathcal{I}}(\boldsymbol{y}\mid\theta)-\int\psi(\boldsymbol{y},\theta)\ \textup{d}\mu^{\ast}(\boldsymbol{y}\mid\theta)\right\rvert\leq\omega_{\psi}(\delta)$ (5.25)
and for any Bayesian Wardrop $\alpha$-equilibrium:
$\left\lvert\int\psi(\boldsymbol{y},\theta)\ \textup{d}\mu^{\ast}(\boldsymbol{y}\mid\theta)-\int\psi(\boldsymbol{y},\theta)\ \textup{d}\widetilde{\mu}(\boldsymbol{y}\mid\theta)\right\rvert\leq\omega_{\psi}(\delta).$ (5.26)
It follows that
$\left\lvert\int\psi(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)-\int\psi(\boldsymbol{y},\theta)\ \textup{d}\widetilde{\mu}(\boldsymbol{y}\mid\theta)\right\rvert\leq\alpha+2\omega_{\psi}(\delta),$ (5.27)
which is smaller than $\varepsilon$ for $\alpha,\delta$ small enough. ∎
Proof of Theorem˜5.5.
The result follows from the fact that the unique equilibrium outcome $\mu^{\ast}$ is an Bayesian Wardrop $\alpha$-equilibrium outcome for any $\alpha>0$. As in the proof of Theorem˜4.1, if $\mu$ has finite support with rational flows, then we can choose $\varepsilon=0$, and the obedient outcome $\breve{\mu}_{\mathcal{I}}$ is the unique Bayesian Wardrop equilibrium outcome $\mu^{\ast}$. ∎
The next theorem shows that in games with a weakly concave potential we can fully implement the ex-ante expected total payoff of any Bayes correlated Wardrop equilibrium, where the total payoff in state $\theta$ given flows $\boldsymbol{y}$ is:
$\operatorname{\mathsf{U}}(\boldsymbol{y},\theta)\coloneqq\sum_{a}y_{a}u_{a}(\boldsymbol{y},\theta).$ (5.28)
Theorem 5.9.
Let $\Gamma$ be a basic game with a concave potential in each state and let $\mu$ be a BCWE. For all $\varepsilon>0$, there exists a direct information structure $\mathcal{I}$ such that for each Bayesian Wardrop equilibrium outcome $\mu^{\ast}$ of $(\Gamma,\mathcal{I})$,
$\left\lvert\sum_{\theta}p(\theta)\int\operatorname{\mathsf{U}}(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)-\sum_{\theta}p(\theta)\int\operatorname{\mathsf{U}}(\boldsymbol{y},\theta)\ \textup{d}\mu^{\ast}(\boldsymbol{y}\mid\theta)\right\rvert\leq\varepsilon.$ (5.29)
As before, if $\mu$ has finite support and rational flows, we can choose $\varepsilon=0$ and we know that the obedient outcome $\breve{\mu}_{\mathcal{I}}$ is a Bayesian Wardrop equilibrium outcome.
Theorem˜5.9 applies to Cournot games. In Example˜3.2, it implies that there exists an information structure (the one described on page 3.2) that implements a BCWE which maximizes total profit in every Bayesian Wardrop equilibrium of the basic game extended with this information structure. However, Theorem˜5.9 does not apply to Example˜3.1, which is a potential game but lacks concavity. In this case, Theorem˜4.1 implies that there exists an information structure (the one described on page 3.4) that implements the BCWE which maximizes the total payoff, in a Bayesian Wardrop equilibrium of the basic game extended with this information structure, but other, less efficient equilibria also exist. [^7] Theorem 5.9 does not apply to games with positive payoff externalities such as those in Li et al., (2023). However, Li et al., (2023) provide an example showing that full implementation may still be possible with complex information structures involving an infinite set of signals.
Proof of Theorem˜5.9.
Assume that $\mu$ is a BCWE with finite support, choose the function $\psi$ as the social payoff function $\operatorname{\mathsf{U}}$, and fix $\alpha>0$. Consider the information structure $\mathcal{I}$ constructed in the proof of Theorem˜4.1 with parameter $\eta>0$, such that the obedient flow profile $\breve{\boldsymbol{y}}_{\mathcal{I}}$ with outcome $\breve{\mu}_{\mathcal{I}}$ is a Bayesian Wardrop $\alpha$-equilibrium and
$\left\lvert\int\operatorname{\mathsf{U}}(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)-\int\operatorname{\mathsf{U}}(\boldsymbol{y},\theta)\ \textup{d}\breve{\mu}_{\mathcal{I}}(\boldsymbol{y}\mid\theta)\right\rvert\leq\alpha,\quad\forall\theta\in\Theta.$ (5.30)
We know that the set of Wardrop equilibria of the associated multi-population game with complete information is the set of maximizers of the potential (Sandholm,, 2001, Proposition 3.1). Because there is a concave potential in each state, the potential defined in Eq.˜5.14 is a concave function and the set of Wardrop equilibria is convex. Also, from Koessler et al., (2025, Proposition 4), all Wardrop equilibria have the same payoff profile. That is, for any Wardrop equilibrium flow vectors $\boldsymbol{y}^{*}=(\boldsymbol{y}^{*}(\boldsymbol{\tau}))_{\boldsymbol{\tau}}$ and $\boldsymbol{y}^{**}=(\boldsymbol{y}^{**}(\boldsymbol{\tau}))_{\boldsymbol{\tau}}$, for every $k,\tau^{k}$,
$U_{a^{*}}^{k,\tau^{k}}(\boldsymbol{y}^{*})=U_{a^{**}}^{k,\tau^{k}}(\boldsymbol{y}^{**}),$ (5.31)
where $a^{*}$ (resp. $a^{**}$) is any action in the support of $\boldsymbol{y}^{*}$ (resp. $\boldsymbol{y}^{**}$) for population $(k,\tau^{k})$. It follows that all Wardrop equilibria have the same total payoff, where the total payoff of an interim flow profile $\widehat{\boldsymbol{y}}$ is,
[…]
From (5.13), the total payoff in the population game is equal to the expected total payoff in the basic game, i.e.,
$\begin{split}\widehat{\operatorname{\mathsf{U}}}(\widehat{\boldsymbol{y}})&=\sum_{k,\tau^{k}}\sum_{a}y_{a}^{k}(\tau^{k})\sum_{\theta,\boldsymbol{\tau}^{-k}}p(\theta)\pi\left(\tau^{k},\boldsymbol{\tau}^{-k}\mid\theta\right)u_{a}(\boldsymbol{y}(\boldsymbol{\tau}),\theta)\\ &=\sum_{\theta}p(\theta)\sum_{\boldsymbol{\tau}}\pi\left(\boldsymbol{\tau}\mid\theta\right)\sum_{a}\sum_{k}y_{a}^{k}(\tau^{k})u_{a}(\boldsymbol{y}(\boldsymbol{\tau}),\theta)\\ &=\sum_{\theta}p(\theta)\sum_{\boldsymbol{\tau}}\pi\left(\boldsymbol{\tau}\mid\theta\right)\sum_{a}y_{a}(\boldsymbol{\tau})u_{a}(\boldsymbol{y}(\boldsymbol{\tau}),\theta)\\ &=\sum_{\theta}p(\theta)\sum_{\boldsymbol{\tau}}\pi\left(\boldsymbol{\tau}\mid\theta\right)\operatorname{\mathsf{U}}(\boldsymbol{y},\theta).\end{split}$ (5.33)
It follows that the expected total payoff is the same in all Bayesian Wardrop equilibrium outcomes of $\Gamma$ with information structure $\mathcal{I}$. From ˜5.7, any Bayesian Wardrop $\alpha$-equilibrium of the game with information structure $\mathcal{I}$, is a $\alpha$-maximizer of the associated potential. To complete the proof, we use the following argument:
Claim 5.10.
For all $\delta>0$, there exists $\alpha>0$ such that for all $\widehat{\boldsymbol{y}}$, $\Phi_{\pi}(\widehat{\boldsymbol{y}})\geq\max\Phi_{\pi}-\alpha$ implies
$\left\lvert\widehat{\operatorname{\mathsf{U}}}(\widehat{\boldsymbol{y}})-\widehat{\operatorname{\mathsf{U}}}(\widehat{\boldsymbol{y}}^{*})\right\rvert\leq\delta,$ (5.34)
where $\widehat{\boldsymbol{y}}^{*}$ is any Bayesian Wardrop equilibrium of $\Gamma$ with information structure $\mathcal{I}$.
Proof of ˜5.10.
By contradiction, suppose that there exists $\delta>0$ such that for all $\alpha=1/n$ with $n\in\mathbb{N}$, there exists $\widehat{\boldsymbol{y}}_{n}$ with $\Phi_{\pi}(\widehat{\boldsymbol{y}}_{n})\geq\max\Phi_{\pi}-1/n$ and $\left\lvert\widehat{\operatorname{\mathsf{U}}}(\widehat{\boldsymbol{y}}_{n})-\widehat{\operatorname{\mathsf{U}}}(\widehat{\boldsymbol{y}}^{*})\right\rvert>\delta$. By compactness, take a subsequence and assume that $\widehat{\boldsymbol{y}}_{n}\to\widehat{\boldsymbol{y}}_{\infty}$ as $n\to\infty$. Letting $n\to\infty$, we get that $\Phi_{\pi}(\widehat{\boldsymbol{y}}_{\infty})\geq\max\Phi_{\pi}$ and $\left\lvert\widehat{\operatorname{\mathsf{U}}}(\widehat{\boldsymbol{y}}_{\infty})-\widehat{\operatorname{\mathsf{U}}}(\widehat{\boldsymbol{y}}^{*})\right\rvert\geq\delta$. This is a contradiction because $\widehat{\boldsymbol{y}}_{\infty}$ would then be a maximizer of the potential, thus a Bayesian Wardrop equilibrium, with a different total payoff. ∎
This concludes the proof of Theorem˜5.9. ∎
6 Optimality of Deterministic BCWE
The following proposition identifies a class of games with binary action sets in which, in order to maximize the designer’s payoff function $\psi$, attention can be restricted to deterministic flows conditional on the state. A BCWE $\mu$ is said to be deterministic if, in each state $\theta\in\Theta$, a single flow profile is implemented with probability one. Hence, a deterministic BCWE can be identified with a mapping $\boldsymbol{y}(\,\cdot\,)\colon\Theta\to\mathcal{Y}$ such that, for all $a,b\in\mathcal{A}$,
$\sum_{\theta\in\Theta}p(\theta)\,y_{a}(\theta)u_{a}(\boldsymbol{y}(\theta),\theta)\geq\sum_{\theta\in\Theta}p(\theta)\,y_{a}(\theta)u_{b}(\boldsymbol{y}(\theta),\theta).$ (6.1)
Under complete information, the notion of deterministic BCWE coincides with that of a Wardrop equilibrium.
Proposition 6.1.
Consider a game where $\mathcal{A}$ contains exactly two actions; for every $a\in\mathcal{A}$ and $\theta\in\Theta$, the mapping $\boldsymbol{y}\mapsto u_{a}(\boldsymbol{y},\theta)$ is convex and the mapping $\boldsymbol{y}\mapsto y_{a}u_{a}(\boldsymbol{y},\theta)$ is concave. Moreover, assume that the designer payoff function $\boldsymbol{y}\mapsto\psi(\boldsymbol{y},\theta)$ is concave for every $\theta\in\Theta$. Then, for every BCWE of $\Gamma$, there exists a deterministic BCWE for which the expected designer payoff is weakly higher.
Proof.
Consider a BCWE outcome $\mu$. Then
$\begin{split}&\forall a,b\in\mathcal{A},\ \sum_{\theta\in\Theta}p(\theta)\int y_{a}u_{a}(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)\geq\sum_{\theta\in\Theta}p(\theta)\int y_{a}u_{b}(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta).\end{split}$ (6.2)
Since there are two actions, we have $y_{b}=1-y_{a}$ for $b\neq a$. The incentive constraint corresponding to recommendation $a$ and deviation $b\neq a$ can therefore be rewritten as
$\sum_{\theta\in\Theta}p(\theta)\int y_{a}u_{a}(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)\geq\sum_{\theta\in\Theta}p(\theta)\int(1-y_{b})u_{b}(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta),$ (6.3)
or equivalently,
$\begin{split}&\sum_{\theta\in\Theta}p(\theta)\int y_{a}u_{a}(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)+\sum_{\theta\in\Theta}p(\theta)\int y_{b}u_{b}(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)\\ &\quad\geq\sum_{\theta\in\Theta}p(\theta)\int u_{b}(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta).\end{split}$ (6.4)
For each state $\theta$, define the average flow profile
$\bar{\boldsymbol{y}}(\theta)\coloneqq\int\boldsymbol{y}\ \textup{d}\mu(\boldsymbol{y}\mid\theta),\quad\text{with components }\bar{y}_{a}(\theta)=\int y_{a}\ \textup{d}\mu(\boldsymbol{y}\mid\theta).$ (6.5)
By concavity of $\boldsymbol{y}\mapsto y_{a}u_{a}(\boldsymbol{y},\theta)$ and convexity of $\boldsymbol{y}\mapsto u_{b}(\boldsymbol{y},\theta)$, Jensen’s inequality yields
$\sum_{\theta\in\Theta}p(\theta)\int y_{a}u_{a}(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)$ $\leq\sum_{\theta\in\Theta}p(\theta)\,\bar{y}_{a}(\theta)u_{a}(\bar{\boldsymbol{y}}(\theta),\theta),$ (6.6)
$\sum_{\theta\in\Theta}p(\theta)\int y_{b}u_{b}(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)$ $\leq\sum_{\theta\in\Theta}p(\theta)\,\bar{y}_{b}(\theta)u_{b}(\bar{\boldsymbol{y}}(\theta),\theta),$ (6.7)
$\sum_{\theta\in\Theta}p(\theta)\int u_{b}(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)$ $\geq\sum_{\theta\in\Theta}p(\theta)\,u_{b}(\bar{\boldsymbol{y}}(\theta),\theta).$ (6.8)
Thus, Eq.˜6.4 implies that $\theta\mapsto\bar{\boldsymbol{y}}(\theta)$ satisfies the deterministic BCWE incentive constraints. Finally, concavity of $\boldsymbol{y}\mapsto\psi(\boldsymbol{y},\theta)$ implies
$\sum_{\theta\in\Theta}p(\theta)\int\psi(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)\leq\sum_{\theta\in\Theta}p(\theta)\psi(\bar{\boldsymbol{y}}(\theta),\theta),$ (6.9)
which completes the proof. ∎
Appendix A List of Symbols and Acronyms
$\mathcal{A}$ action set, defined in Definition˜2.1
$a$ action
$b$ action
$\operatorname{\mathsf{BWE_{\varepsilon}}}(\mathcal{I})$ set of Bayesian Wardrop $\varepsilon$-equilibrium outcomes of $\Gamma$ with information structure $\mathcal{I}$
$c_{e}$ cost of resource $e$
$C^{h}$ class of functions with a continuous $h$-th derivative
$e$ resource in a congestion game
$\mathcal{E}$ resource set in a congestion game
$\mathcal{I}$ $\mathcal{I}=\left(\boldsymbol{\gamma},\mathcal{T},\pi\right)$, information structure, defined in Definition˜2.3
$k$ population
$\mathcal{K}$ population set
$L$ Lipschitz constant
$p$ probability distribution on $\Theta$, defined in Definition˜2.1
$\operatorname{\mathsf{P}}(\tau^{k})$ $\sum_{\theta\in\Theta}\sum_{\boldsymbol{\tau}^{-k}\in\mathcal{T}^{-k}}p(\theta)\pi(\tau^{k},\boldsymbol{\tau}^{-k}\mid\theta)$, total probability of type $\tau^{k}$
$\operatorname{\mathsf{Q}}^{k}(a)$ $\sum_{\theta\in\Theta}\sum_{\boldsymbol{y}\in\mathcal{Y}^{\ast}}p(\theta)\mu(\boldsymbol{y}\mid\theta)y_{a}^{k}$, defined in (4.6)
$Q(\boldsymbol{y})$ aggregate quantity in an aggregative game, defined before (5.7)
$\mathcal{T}$ $\mathcal{T}=\times_{k\in\mathcal{K}}\mathcal{T}^{k}$, defined in Definition˜2.3
$\mathcal{T}^{k}$ set of types for population $k$, defined in Definition˜2.3
$\mathcal{T}^{-k}$ $\times_{j\neq k}\mathcal{T}^{j}$
$u_{a}$ payoff of action $a$, defined in Definition˜2.1
$u_{a}^{k,\tau^{k}}$ payoff functions of population $(k,\tau^{k})$, defined in (2.4)
$\boldsymbol{u}$ $(u_{a})_{a\in\mathcal{A}}$, defined in Definition˜2.1
$(u_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}}))_{k,\tau^{k},a}$ interim payoff profile
$U_{a}^{k,\tau^{k}}$ payoff function of population $(k,\tau^{k})$, defined in (5.13)
$\operatorname{\mathsf{U}}(\boldsymbol{y},\theta)$ $\sum_{a}y_{a}u_{a}(\boldsymbol{y},\theta)$, defined in (5.28)
$\operatorname{\mathsf{WE}}(\Gamma)$ Wardrop equilibria of $\Gamma$
$x_{e}$ load on resource $e$, defined after (5.2)
$y_{a}^{k}(\tau^{k})$ flow on action $a$ of population $k$, when type $\tau^{k}$ is observed
$y_{a}(\boldsymbol{\tau})$ $\sum_{k\in\mathcal{K}}y_{a}^{k}(\tau^{k})$, interim total flow on action $a$
$\breve{\boldsymbol{y}}_{\mathcal{I},a}^{k}(\tau^{k})$ $\gamma^{k}\cdot\operatorname{\mathds{1}}\left\{\tau^{k}=a\right\}$
$\boldsymbol{y}$ flow profile
$\boldsymbol{y}^{\ast}$ equilibrium flow profile
$\boldsymbol{y}^{k}(\tau^{k})$ $\left(y_{a}^{k}(\tau^{k})\right)_{a\in\mathcal{A}}\in\Delta_{\gamma^{k}}(\mathcal{A})$, interim flow profile of population $k$ given type $\tau^{k}$
$\boldsymbol{y}(\boldsymbol{\tau})$ $\left(y_{a}(\boldsymbol{\tau})\right)_{a\in\mathcal{A}}\in\Delta(\mathcal{A})$, interim total flow profile given $\boldsymbol{\tau}$
$\widehat{\boldsymbol{y}}$ $\left(\boldsymbol{y}^{k}(\tau^{k})\right)_{k,\tau^{k}}\in\bigtimes_{k\in\mathcal{K}}\left[\bigtimes_{\tau^{k}\in\mathcal{T}^{k}}\Delta_{\gamma^{k}}(\mathcal{A})\right]$, interim flow profile, defined in (2.3)
$\breve{\boldsymbol{y}}_{\mathcal{I}}$ obedient interim flow profile
$\mathcal{Y}$ $\Delta_{\gamma}(\mathcal{A})$, set of flow profiles, defined before Definition˜2.1
$\gamma^{k}$ size of population $k$, defined in Definition˜2.3
$\boldsymbol{\gamma}$ vector of population sizes, defined in Definition˜2.3
$\Gamma$ $\left(\mathcal{A},\Theta,p,\boldsymbol{u}\right)$, basic game, defined in Definition˜2.1
$\delta_{a}$ Dirac measure on $a$
$\Delta(A)$ $\Delta_{1}(A)$, simplex of probability measures on AA
$\Delta_{\gamma}(A)$ $\left\{\boldsymbol{y}\in\mathbb{R}^{A}\colon\forall a\in A,y_{a}\geq 0,\sum_{a\in A}y_{a}=\gamma\right\}$, defined in (1.1)
$\theta$ state
$\Theta$ state space, defined in Definition˜2.1
$\mu$ outcome of the game, defined after Definition˜2.1
$\mu^{\ast}$ Bayesian Wardrop equilibrium outcome
$\breve{\mu}_{\mathcal{I}}$ outcome associated to the obedient interim flow profile
$\pi$ mapping from $\Theta$ to $\Delta(\mathcal{T})$, defined in Definition˜2.3
$\tau^{k}$ possible type for population $k$
$\boldsymbol{\tau}$ $(\tau^{k})_{k\in\mathcal{K}}$ type profile
$\Phi_{\theta}$ potential, defined in Definition˜5.1
$\psi$ designer welfare function, first used in Theorem˜4.1
$\omega(\,\cdot\,)$ modulus of continuity
$\operatorname{\mathds{1}}$ indicator function
BCWE Bayes correlated Wardrop equilibrium, defined in Definition˜2.2
BWE Bayesian Wardrop equilibrium, defined after (2.6)
[…]